### 11 Mar 2026 v1

# The Kobayashi-Hitchin correspondence for nef and big classes

Abstract
[…]
1 Introduction
The Kobayashi-Hitchin correspondence asserts that a holomorphic vector bundle $E$ over a compact complex manifold $X$ is slope polystable with respect to a Kähler class $\{\omega\}$ on $X$ if and only if $E$ admits an $\omega$-Hermitian-Yang-Mills metric. If $\omega$ is a smooth Kähler metric, this correspondence was proved by Donaldson [Don85] and Uhlenbeck-Yau [UY86] for holomorphic vector bundles on compact Kähler manifolds, by Bando-Siu [BS94] for reflexive sheaves on compact Kähler manifolds by introducing the notion of admissible Hermitian-Yang-Mills metrics, and by Xuemiao Chen [Chen25] for reflexive sheaves on compact normal Kähler varieties. More recently, the correspondence has been extended to compact klt Kähler varieties endowed with orbifold Kähler metrics motivated by the study of the equality cases of the Bogomolov-Gieseker inequality and the Miyaoka-Yau inequality of klt varieties [Faulk22], [FO25] (refer to the introduction of [IJZ25] for more detail). There are also results when Kähler metrics have conical singularities [Li00], [JL25].
In [Jin25], the author defined the $\langle\alpha^{n-1}\rangle$-slope stability for a big cohomology class $\alpha$ and established the Kobayashi-Hitchin correspondence on a compact complex manifold equipped with a big line bundle $L$, under the assumption that $L$ admits an ample model. Recent developments in the study of the Yau-Tian-Donaldson conjecture for cscK metrics have highlighted the importance of big line bundles that do not necessarily admit ample models [BJ25]. Any closed positive $(1,1)$-current with minimal singularities representing a big class is not positive definite and generally has singularities that cannot be described explicitly. Consequently, the development of analytic techniques adapted to such positive currents has become increasingly important.
[…]
Our approach is based on the Uhlenbeck compactness theorem for Yang-Mills connections, which is inspired by the approach of [Chen25]. We are also inspired by the arguments in [KO25].
1.1 Main results
1.1.1 Definitions
Our main result of this paper is Theorem 1.4. To formulate Theorem 1.4, we firstly introduce the notion of adapted closed positive $(1,1)$-currents, inspired by the notion of tame approximations of [CCHTST25] for closed positive $(1,1)$-currents with bounded potentials (refer to Definition 2.18 for singular settings).
Definition 1.1 (=Definition 2.15, Definition 2.18).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Then, we say that a closed positive $(1,1)$-current $T$ lying in $\alpha$ is adapted if it satisfies the following conditions: Let $\pi:Y\to X$ be a sequence of blow-ups so that $D=E_{nK}(\pi^{*}\alpha)$, the non-Kähler locus of $\pi^{*}\alpha$, is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. Then,
  1. $(1)$
    $T$ is smooth Kähler on ${\rm{Amp}}(\alpha)$, the ample locus of $\alpha$.
  2. $(2)$
    $\pi^{*}T\geq|s_{D}|^{2m}_{h_{D}}\omega_{Y}$ for some $m\in\mathbb{Z}_{>0}$, where $\omega_{Y}$ is a smooth Kähler metric on $Y$.
  3. $(3)$
    There exists a sequence of smooth Kähler metrics $\omega_{i}$ lying in $\pi^{*}\alpha+(1/i)\{\omega_{Y}\}$ such that $\omega_{i}$ locally smoothly converges to $\pi^{*}T$ on $Y\setminus D={\rm{Amp}}(\pi^{*}\alpha)$ and
    $\int_{Y}\left(\frac{\omega_{i}^{n}}{\omega_{Y}^{n}}\right)^{p}\omega_{Y}^{n}\leq C$ (1.1)
    for some constant $C>0$ and $p>1$ which are both independent of $i$. We say the above $\{\omega_{i}\}_{i}$ as an adapted approximation of $T$.
We will see that any nef and big class contains an adapted current (Proposition 2.16). One of the important examples of adapted currents is a singular Kähler-Einstein metric on a compact normal Kähler variety with log terminal singularities (Theorem 2.20). We next introduce the notion of adapted Hermitian-Yang-Mills metrics as admissible Hermitian-Yang-Mills metrics with additional two conditions, called as adapted conditions (refer to Definition 2.22 for the definition in singular settings):
Definition 1.2 (=Definition 2.21, Definition 2.22).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$, the ample locus of $\alpha$. Let $E$ be a holomorphic vector bundle on $X$ with a smooth hermitian metric $h_{0}$ on $E$. Then, a smooth hermitian metric $h=h_{0}\Psi^{2}$ on $E|_{{\rm{Amp}}(\alpha)}$ is said to be a $T$-adapted Hermitian-Yang-Mills metric if it satisfies the following four conditions:
  1. $(1)$
    $\sqrt{-1}\Lambda_{T}F_{h}=\lambda{\rm{Id}}$ on ${\rm{Amp}}(\alpha)$ for some $\lambda\in\mathbb{R}$.
  2. $(2)$
    $\int_{{\rm{Amp}}(\alpha)}|F_{h}|^{2}_{h,T}T^{n}<\infty$.
Let $\pi:Y\to X$ be a sequence of blow-ups so that $D=E_{nK}(\pi^{*}\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. Then
  1. $(3)$
    $\sup_{Y\setminus D}|s_{D}^{k}\pi^{*}\Psi|_{\pi^{*}h_{0}\otimes\pi^{*}h_{0}^{*}\otimes h_{D}^{k}}<\infty$ for some $k\in\mathbb{Z}_{>0}$.
  2. $(4)$
    $\int_{Y\setminus D}|s_{D}^{l}\partial^{\nabla_{\pi^{*}h_{0}}}(\pi^{*}\Psi)|^{2}_{\pi^{*}h_{0}\otimes\pi^{*}h_{0}^{*}\otimes h_{D}^{l},T}(\pi^{*}T)^{n}<\infty$ for some $l\in\mathbb{Z}_{>0}$.
We call $\lambda$ in (1) the $T$-Hermitian-Yang-Mills constant of $h$ (refer to Corollary 3.24 for the explicit description of $\lambda$).
We will see more precise estimates on $T$-adapted Hermitian-Yang-Mills metrics for adapted currents $T$ (refer to Corollary 1.9). In [UY86], Uhlenbeck-Yau introduced the notion of weak holomorphic subbundles which is a Sobolev theoretic analog of orthogonal projections to holomorphic subbundles. It plays an important role in the proof of [UY86]. In this paper, we generalize their weak holomorphic subbundles to the case that Kähler metrics on base manifolds are replaced by closed positive $(1,1)$-currents.
Definition 1.3 (=Definition 3.11).
Let $X$ be a compact Kähler manifold and $T$ be a closed positive $(1,1)$-current on $X$ which is smooth Kähler on $X\setminus D$ where $D$ is a complex analytic subset in $X$. Let $(E,h_{0})$ be a holomorphic hermitian vector bundle on $X$. Then, a $T$-weak holomorphic subbundle of $E$ is an $h_{0}$-hermitian endomorphism $\pi$ of $E$ defined almost everywhere on $X\setminus D$ such that $\pi^{2}=\pi$, $(1-\pi)\circ\bar{\partial}^{E}\pi=0$ and
$\int_{X\setminus D}|\pi|_{h_{0}}^{2}T^{n}+\int_{X\setminus D}|\nabla_{h_{0}}\pi|^{2}_{h_{0},T}T^{n}<\infty.$
By [UY86], the image of a $T$-weak holomorphic subbundle $\pi$ uniquely defines a coherent subsheaf $\mathcal{F}$ of $E|_{X\setminus D}$. We denote $\mathcal{F}$ by ${\rm im}(\pi)$.
We will see that a $T$-weak holomorphic subbundle defines a torsion free coherent subsheaf $\mathcal{F}\subset E$ defined all over $X$ and it satisfies the Chern-Weil formula (refer to Proposition 1.5).
1.1.2 The Kobayashi-Hitchin correspondence
The main result of this paper is the following:
Theorem 1.4 (=Theorem 3.17, Theorem 3.28, Theorem 3.29).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T$ be an adapted closed positive $(1,1)$-current in $\alpha$. Let $E$ be a holomorphic vector bundle over $X$. Then, the following conditions are equivalent:
  1. $(1)$
    $E$ is $\alpha^{n-1}$-slope polystable (refer to Definition 2.10).
  2. $(2)$
    $E$ admits a $T$-adapted Hermitian-Yang-Mills metric.
Furthermore, if $E$ admits a $T$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
We remark that the proof of (2) $\Rightarrow$ (1) is also nontrivial in this setting and the proof is presented in §3.4. In the proof of (1) $\Rightarrow$ (2) in Theorem 1.4, the following generalization of the Chern-Weil formula to $T$-weak holomorphic subbundles plays an important role:
Proposition 1.5 (=Proposition 3.12).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T$ be a closed positive $(1,1)$-current with minimal singularities in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$. Let $(E,h_{0})$ be a holomorphic hermitian vector bundle on $X$. Let $\pi$ be a $T$-weak holomorphic subbundle of $E$. Then the following holds.
  1. $(1)$
    There exists a torsion free coherent subsheaf $\mathcal{F}\subset E$ on $X$ such that $\mathcal{F}|_{{\rm{Amp}}(\alpha)}={\rm im}(\pi)$.
  2. $(2)$
    There exists an analytic subset $W\subset X$ such that
    $\int_{X}c_{1}(\mathcal{F})\wedge\alpha^{n-1}=\int_{X\setminus W}\left({\rm{Tr}}(\pi\sqrt{-1}F_{h_{0}})-{\rm{Tr}}(\bar{\partial}\pi\wedge(\bar{\partial}\pi)^{*_{h_{0}}})\right)\wedge\langle T^{n-1}\rangle.$
The following theorem is a direct consequence of Theorem 1.4 via resolution of singularities (refer to Remark 3.1).
Corollary 1.6 (refer to Lemma 2.12, Lemma 2.19 and Lemma 2.23, Theorem 1.4).
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\mathcal{E}$ be a reflexive sheaf on $X$. Then the following conditions are equivalent.
  1. $(1)$
    $\mathcal{E}$ is $\alpha^{n-1}$-slope polystable (refer to Definition 2.11).
  2. $(2)$
    $\mathcal{E}$ admits a $T$-adapted Hermitian-Yang-Mills metric (refer to Definition 2.22).
Furthermore, if $\mathcal{E}$ admits a $T$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
[…]
Theorem 1.7 (=Theorem 3.32).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\mathcal{E}$ be a torsion free sheaf on $X$. Suppose that $\mathcal{E}$ is $\alpha^{n-1}$-slope semistable. Then the following holds.
  1. $(1)$
[…]
  2. $(2)$
[…]
It is an interesting problem whether the $\omega_{i}$-Hermitian-Yang-Mills flow converges the $T$-adapted Hermitian-Yang-Mills metric where $\omega_{i}$ is an adapted approximation of $T$ (refer to Conjecture 3.33).
By [CCHTST25], it has proved that a singular Kähler-Einstein metric on a compact klt variety is an adapted current. Hence we obtain the following result as a direct consequence of Corollary 1.6. Following Corollary 1.8 itself is a new result to the best of our knowledge.
Corollary 1.8 (see also [Jin25-2], [Jin25-3]).
Let $X$ be a compact normal Kähler variety with log terminal singularities. Suppose that $X$ admits a singular Kähler-Einstein metric $\omega$. Let $\mathcal{E}$ be a reflexive sheaf on $X$. Then the following conditions are equivalent:
  1. $(1)$
    $\mathcal{E}$ is $\{\omega\}^{n-1}$-slope polystable.
  2. $(2)$
    $\mathcal{E}$ admits an $\omega$-adapted Hermitian-Yang-Mills metric.
Furthermore, if $\mathcal{E}$ admits an $\omega$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
1.1.3 A priori estimates and applications
As a consequence of the proof of Theorem 1.4, we obtain the following a priori estimates on $T$-adapted HYM metrics. We remark that we can assume that the non-Kähler locus is a snc divisor without loss of generalities (refer to Remark 3.1).
Corollary 1.9 (=Proposition 3.15, Corollary 3.30).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$ such that $D=E_{nK}(\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. Let $T$ be an adapted closed positive $(1,1)$-current lying in $\alpha$. Let $m\in\mathbb{Z}_{\geq 0}$ be a nonnegative integer such that $T\geq|s_{D}|^{2m}\omega_{0}$. Assume that a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric $h=h_{0}\Psi^{2}$ with $\det\Psi=1$. Then the following holds:
  1. $(1)$
    $\sup_{X\setminus D}|s_{D}\Psi|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty$.
  2. $(2)$
    $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi|_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},T}T^{n}<\infty$.
  3. $(3)$
    We denote by $\Psi=e^{S}$ for some $h_{0}$-hermitian endomorphism $S$. Then the Lelong-type number of $S$ relative to $\log|s_{D}|_{h_{D}}^{2}$ equals to 0 in the following sense: Let $\varepsilon_{k}>0$ be a sequence of constants so that $\varepsilon_{k}\searrow 0$. We denote by $D_{\varepsilon_{k}}$ the $\varepsilon_{k}$-neighborhood of $D$ with respect to $\omega_{0}$. Then
    $\inf_{k}\sup_{D_{\varepsilon_{k}}}\frac{|S|_{h_{0}}}{-\log|s_{D}|^{2}}=0.$
If $\omega$ is a Kähler class on a compact Kähler manifold $X$, then it is well-known that any holomorphic endomorphism of an $\omega^{n-1}$-slope stable vector bundle is constant (e.g.[Kob87, Chapter 5, Corollary 7.14]). The following is a generalization of this theorem to the nef and big class-setting. Proposition 1.10 is used in the proof of the uniqueness of $T$-adapted HYM metrics (see §1.2 below) and is of independent interest. We again remark that we can assume that the non-Kähler locus is a snc divisor without loss of generalities (refer to Remark 3.1).
Proposition 1.10 (=Lemma 3.25).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$ with $D=E_{nK}(\alpha)$ is a snc divisor. Let $E_{1}$ and $E_{2}$ be $\alpha^{n-1}$-slope stable holomorphic vector bundles on $X$ with the same rank ${\rm{rk}}E_{1}={\rm{rk}}E_{2}$ and the same $\alpha^{n-1}$-slope $\mu_{\alpha}(E_{1})=\mu_{\alpha}(E_{2})$. Let $g:E_{1}|_{X\setminus D}\to E_{2}|_{X\setminus D}$ be a nontrivial holomorphic morphism satisfying
$\sup_{X\setminus D}|s_{D}^{l}g|_{h_{D}\otimes h_{2}\otimes h_{1}^{*}}<\infty,$
for some $l\in\mathbb{Z}_{>0}$, where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ and $h_{i}$ is a smooth hermitian metric on $E_{i}$ ($i=1,2$). Then $g$ is an isomorphism. In particular, if $E=E_{1}=E_{2}$, then any nontrivial holomorphic endomorphism $g$ of $E$ is a constant multiplication of an identity morphism, that is, $g=a{\rm{id}}_{E}$ for some $a\in\mathbb{C}\setminus\{0\}$.
Since the direct sum, the tensor product and the wedge product of $T$-adapted HYM metrics are again $T$-adapted HYM, we obtain the following result as a consequence of Corollary 1.6.
Corollary 1.11.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Then, for any $\alpha^{n-1}$-slope polystable reflexive sheaves $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, their direct sum $\mathcal{E}_{1}\oplus\mathcal{E}_{2}$, tensor product $\mathcal{E}_{1}\otimes\mathcal{E}_{2}$ and wedge product $\mathcal{E}_{1}\wedge\mathcal{E}_{2}$ are $\alpha^{n-1}$-slope polystable.
As an application of Corollary 1.6, we obtain the following theorem.
Theorem 1.12.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\mathcal{E}$ be an $\alpha^{n-1}$-slope polystable reflexive sheaf on $X$. Then $\mathcal{E}$ satisfies the following Bogomolov-Gieseker inequality:
$\left(2rc_{2}(\mathcal{E})-rc_{1}(\mathcal{E})^{2}\right)\cdot\alpha^{n-2}\geq 0.$ (1.2)
[…]
By the proof of Theorem 1.4, the proof of Theorem 1.12 is the same with [Chen25, §4.1.1]. The Bogomolov-Gieseker inequality (1.2) was proved by [IJZ25] for big, not necessarily nef, cohomology classes by replacing $\alpha^{n-2}$ to the positive product $\langle\alpha^{n-2}\rangle$. The projectively flatness in the equality of the Bogomolov-Gieseker inequality as above Theorem 1.12 is even new if $X$ is a projective manifold and $\alpha$ is the 1st Chern class of a nef and big line bundle.
1.1.4 Further discussions
Theorem 1.12 defines the following map:
[…]
If a nef and big class $\alpha$ is a Kähler class, then above map is bijective and it is the non-abelian Hodge correspondence in the case that a Higgs structure is trivial (refer to [ZZZ26] for more detailed and precise statement). Then it is a natural question whether the above map is bijective in this nef and big class-setting. The identification of the image of the above map will be important.
Next question concerns the precise asymptotic behaviour of adapted HYM metrics. It is well-known that the Lelong number of any closed positive $(1,1)$-current with minimal singularities $T$ in a nef and big class equals to 0 at each points in $X$ (see Proposition 2.8). Since $\log|s_{D}|$ is greater than $\log|z-x|$ around $x\in X$ where $z=(z_{1},\ldots,z_{n})$ is a holomorphic coordinate at $x$, Corollary 1.9 (3) does not ensures that $T$-adapted HYM metrics are equisingular with potentials of $T$. Hence the following question naturally arises:
Conjecture 1.13.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T=\theta+dd^{c}\varphi$ be an adapted closed positive $(1,1)$-current with minimal singularities in $\alpha$ where $\theta$ is a smooth representative in $\alpha$. Let $E$ be a holomorphic vector bundle on $E$ with a smooth hermitian metric $h_{0}$. Assume that $E$ admits a $T$-adapted HYM metric $h=h_{0}e^{2S}$ with an $h_{0}$-hermitian endomorphism $S$. Then $S$ is equisingular with $\varphi$, that is, $|S|_{h_{0}\otimes h_{0}^{*}}+\varphi\in L^{\infty}(X)$.
1.2 Outline of the proof of Theorem 1.4
We use notations in Theorem 1.4. Since $\alpha^{n-1}$-slope stability of a holomorphic vector bundle $E$ is invariant under modifications along the non-Kähler locus $E_{nK}(\alpha)$ (Lemma 2.12) and the pull-back of an adapted current along such modifications is again adapted, we can assume that $D=E_{nK}(\alpha)$ is a snc divisor in $X$ and $T$ admits an adapted approximation with a smooth Kähler metric $\omega_{i}\in\alpha+(1/i)\omega_{0}$ on $X$. There exists a constant $k>0$ such that $|s_{D}|^{k}\langle T^{n}\rangle\leq C\omega_{0}^{n}$, since the density function of $\langle T^{n}\rangle$ with respect to $\omega_{0}^{n}$ lies in $L^{p}(X,\omega_{0}^{n})$ with $p>1$ by definition of adapted currrents.
[…]
$\sup_{X\setminus D}|s_{D}\Psi_{\infty}^{\pm 1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty,$ (1.3)
where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$ (refer to Lemma 3.7). In fact, such $\Psi_{\infty}$ defines a hermitian metric $h_{0}\Psi_{\infty}^{2}$ on $E|_{X\setminus D}$ and it gives a $T$-adapted HYM metric (see also Proposition 3.17). In the construction of $\Psi_{\infty}^{\pm 1}$ satisfying (1.3), the mean value type inequality of Guo-Phong-Sturm [GPS24] (Lemma 3.6) plays an important role. In fact, we can easily shows
$\Delta_{\omega_{i}}\left(\log|s_{D}\Psi_{i}^{\pm 1}|^{2}_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}+a\right)\geq-C$
on $\Omega_{a}:=\{\log|s_{D}\Psi_{i}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}+a>0\}$, where $a>0$ is a small constant such that $\Omega_{a}\neq\phi$ for all $i$. Let us denote by $\Psi_{i}=e^{s_{i}}$ for some $h_{0}$-hermitian endomorphism $s_{i}$ of $(E,h_{0})$ with ${\rm{Tr}}(s_{i})=0$. Then the Guo-Phong-Strum’s mean value type inequality shows that
$\sup_{X}(\log|s_{D}|^{2}+|s_{i}|)\leq\log|s_{D}\Psi_{i}^{\pm 1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C\left(1+\int_{X}(-\log|s_{D}|+|s_{i}|)\omega_{i}^{n}\right)$ (1.4)
holds with universal constant $C>0$ (Lemma 3.8). Then, we can prove that if the $L^{1}$ norm in LHS in (1.4), denoted by $l_{i}$, diverges to $\infty$ in $i\to\infty$, then the normalized morphism $u_{i}:=l_{i}^{-1}s_{i}$ converges to a nontrivial limit $u_{\infty}$ with ${\rm{rk}}(u_{\infty})<{\rm{rk}}E$ (Lemma 3.9). Then, as the argument in [Sim88, Section 5], we can see that the projections to subspaces of the kernel of $u_{\infty}$ define $T$-weak holomorphic subbundles of $E$. Then, if we use the Chern-Weil formula for $T$-weak holomorphic subbundles (Proposition 1.5), we can see that some of the projections produce a destabilizing subsheaf of $E$ (Lemma 3.9). It contradicts to the $\alpha^{n-1}$-slope stability of $E$. Hence the $L^{1}$-norm of (1.4) is uniformly bounded and then $\Psi_{i}$ and $\Psi_{i}^{-1}$ converge to nontrivial morphisms $\Psi_{\infty}$ and $\Phi_{\infty}$ respectively satisfying (1.3) (Lemma 3.7 and Lemma 3.8). Then we can see that $\Phi_{\infty}=\Psi_{\infty}^{-1}$ and $\Psi_{\infty}$ gives a $T$-adapted HYM metric $h_{0}\Psi_{\infty}^{2}$ on $E$ (Theorem 3.17).
(2) $\Rightarrow$ (1): In contrast to the smooth case, the proof of this direction is also nontrivial. We assume that $E$ admits two $T$-adapted HYM metrics $h_{1}$ and $h_{2}$. [1st step] We firstly prove that the Chern connection of $h_{1}$ coincides with that of $h_{2}$: $\nabla_{h_{1}}=\nabla_{h_{2}}$ (Theorem 3.23). In this first step, the result in (1) $\Rightarrow$ (2) plays an important role. We denote by $h_{2}=h_{1}g$ for some positive definite $h_{1}$-hermitian endomorphism $g$. Since each $h_{i}$ is $T$-adapted, we have
$\sup_{X\setminus D}|s_{D}^{l}g|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty$ (1.5)
for some $l\in\mathbb{Z}_{>0}$ where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. The first technical lemma is to show that $g$ is holomorphic, that is, the following holds (Claim 3.27)
$\bar{\partial}^{E\otimes E^{*}}g=0$ (1.6)
on $X\setminus D$. [2nd step] Then, by Proposition 1.10, we obtain that a $T$-adapted HYM metric on an $\alpha^{n-1}$-slope stable vector bundle is unique (Lemma 3.26). [3rd step] Then, by applying the existence of $T$-adapted HYM metrics on $\alpha^{n-1}$-slope stable sheaves and 2nd step, we can prove that a holomorphic vector bundle $E$ admitting a $T$-adapted HYM metric is $\alpha^{n-1}$-slope polystable (Theorem 3.28). [4th step] Let $E|_{X\setminus D}=(\bigoplus_{i=1}^{k}\mathcal{E}_{i})|_{X\setminus D}$ be the decomposition into $\alpha^{n-1}$-slope stable torsion free sheaves of same $\alpha^{n-1}$-slope. Then we can prove that, for $j=1,2$, $h_{j}=\oplus_{i=1}^{k}(h_{j}|_{\mathcal{E}_{i}})$ and each $h_{j}|_{\mathcal{E}_{i}}$ defines a $T$-adapted HYM metric on $\mathcal{E}_{i}$ (Theorem 3.29). Hence, by the uniqueness result for $\alpha^{n-1}$-slope stable sheaves, we can see that $h_{1}=h_{2}$.
[…]
Acknowledgment
The author would like to thank his supervisor Ryushi Goto for his helpful comments and nice discussions. The author also thanks Prof. Yoshinori Hashimoto for suggesting the Uhlenbeck compactness to solve the problem and discussions. The author would like to thank Prof. Yuji Odaka for introducing him to the important paper [CCHTST25]. The author is grateful to Prof. Yuta Watanabe and Rei Murakami for fruitful discussions and important comments.
2 Preliminaries and Definitions
2.1 Nonpluripolar product
2.1.1 Positivities of cohomology classes
Let $X$ be a compact Kähler manifold with a smooth Kähler metric $\omega$. In this paper, we will focus on the positivities of cohomology classes in the following sense.
Definition 2.1.
  1. $(1)$
    A cohomology class $\alpha\in H^{1,1}(X,\mathbb{R})$ is pseudo-effective if $\alpha$ is represented by a closed positive $(1,1)$-current.
  2. $(2)$
    A cohomology class $\alpha\in H^{1,1}(X,\mathbb{R})$ is big if $\alpha$ is represented by a Kähler current $T$, that is, there exists $\varepsilon>0$ such that $T\geq\varepsilon\omega$ in the sense of currents.
  3. $(3)$
    A cohomology class $\alpha$ is nef if for any $\varepsilon>0$ there exists a smooth $(1,1)$-form $\alpha_{\varepsilon}$ in $\alpha$ such that $\alpha_{\varepsilon}\geq-\varepsilon\omega$.
It is well-known that a nef cohomology class $\alpha$ is big if and only if its volume $\alpha^{n}$ is positive [DP04].
Definition 2.2 ([Bou04]).
Let $\alpha$ be a pseudo-effective class on $X$. Then, the non-Kähler locus of $\alpha$ is defined as
[…]
The ample locus of $\alpha$ is the complement of the non-Kähler locus of $\alpha$, that is, ${\rm{Amp}}(\alpha):=X\setminus E_{nK}(\alpha)$.
In singular settings, we define the non-Kähler locus in the following way:
Definition 2.3.
Let $X$ be a compact normal variety in Fujiki’s class $\mathcal{C}$ and $\alpha\in H^{1,1}_{BC}(X)$ be a pseudo-effective class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is smooth Kähler. Then, we define the non-Kähler locus of $\alpha$ by
$E_{nK}(\alpha):=\pi(E_{nK}(\pi^{*}\alpha)).$
The ample locus of $\alpha$ is defined by ${\rm{Amp}}(\alpha):=X\setminus E_{nK}(\alpha)$.
Remark 2.4.
In the above Definition 2.3, we can see that $X_{{\rm{sing}}}\subset E_{nK}(\alpha)$.
Boucksom [Bou04, Theorem 3.17] proved that the non-Kähler locus of big class is an analytic subset of $X$. By Demailly’s regularization, the ample locus of any big class is non-empty. Furthermore, a big class $\alpha$ is Kähler if and only if ${\rm{Amp}}(\alpha)=X$. The following result by Collins-Tosatti [CT15] plays an important role:
Theorem 2.5 ([CT15]).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class. Let $D$ be a divisor contained in $E_{nK}(\alpha)$ the non-Kähler locus of $\alpha$. Then the following holds:
$\alpha^{n-1}[D]=0.$
2.1.2 Nonpluripolar product
In this subsection, we review the general definition of nonpluripolar product of closed positive $(1,1)$-currents following [BEGZ10] (refer to [IJZ25] for definition in singular settings). Let $T$ be a closed positive $(1,1)$-current in a pseudo-effective class $\alpha$. Let $\varphi$ be a plurisubharmonic function such that $T=dd^{c}\varphi$ locally. Bedford-Taylor [BT82] proved that if $\varphi$ is locally bounded, then we can define the wedge product $T^{p}$ inductively:
$\langle(dd^{c}\varphi)^{p},\eta\rangle:=\langle\varphi(dd^{c}\varphi)^{p-1},dd^{c}\eta\rangle$
where $\eta$ is a compactly supported smooth form and $\langle\cdot,\cdot\rangle$ here is the natural pairing between currents and smooth forms. Then the nonpluripolar product of positive currents in general are defined as the limit of the Bedford-Taylor’s wedge product on the locally bounded locus. More precisely, the local nonpluripolar product of closed positive $(1,1)$-current $T=dd^{c}\varphi$ is defined as
$\langle(dd^{c}\varphi)^{p}\rangle:=\lim_{k\to\infty}1_{\{\varphi>-k\}}(dd^{c}\max{\{\varphi,-k\}})^{p},$
where the limit is the weak limit. We say that the nonpluripolar product of $T$ is well defined if it has locally finite mass. On compact complex, not necessarily Kähler, manifolds, it is not known whether the nonpluripolar product of any closed positive $(1,1)$-currents is well defined or not. However, the following was proved by [BEGZ10].
Proposition 2.6 ([BEGZ10], Proposition 1.6, Theorem 1.8).
Let $X$ be a compact Kähler manifold and $T$ be a closed positive $(1,1)$-current in a pseudo-effective class $\alpha$. Then the nonpluripolar product of $T$, denoted by $\langle T^{p}\rangle$, is always globally well-defined and it is a closed positive $(p,p)$-current on $X$.
Next we introduce a simple description of nonpluripolar products in a special setting which is used in this paper. Let $T_{i}$ be a closed positive $(1,1)$-current on $X$ ($i=1,2$) contained in a pseudo-effective cohomology class $\alpha$. We fix a smooth representative $\theta$ of $\alpha$. Then, there exists a $\theta$-plurisubharmonic function $\varphi_{i}$ on $X$, called as a potential of $T_{i}$, such that $T_{i}=\theta+dd^{c}\varphi_{i}$. Then we say that $T_{1}$ is less singular than $T_{2}$ if there is a bounded function $f$ on $X$ such that $\varphi_{2}\leq\varphi_{1}+f$ holds. We say a closed positive $(1,1)$-current $T$ in a pseudo-effective class $\alpha$ has minimal singularities if $T$ is less singular than any other closed positive $(1,1)$-current in $\alpha$. If $\alpha$ is a big class, then it contains a Kähler current on $X$ which is smooth Kähler on the ample locus ${\rm{Amp}}(\alpha)$ ([Bou04, Theorem 3.17]). Hence, any potential of a closed positive $(1,1)$-current $T$ with minimal singularities in a big class $\alpha$ is locally bounded on ${\rm{Amp}}(\alpha)$. Since the nonpluripolar products put no mass on any pluripolar subset, we can see that if $T$ has minimal singularities in a big class $\alpha$, then
$\langle T^{p}\rangle=1_{{\rm{Amp}}(\alpha)}(T|_{{\rm{Amp}}(\alpha)})^{p}$
where the RHS above is the zero extension from ${\rm{Amp}}(\alpha)$ to $X$ of a closed positive $(p,p)$-current $(T|_{{\rm{Amp}}(\alpha)})^{p}$ on ${\rm{Amp}}(\alpha)$.
Definition 2.7 ([BEGZ10], Definition 1.17).
Let $\alpha$ be a pseudo-effective class on a compact Kähler manifold $X$. Then, the positive product of $\alpha$ is defined as the cohomology class of the nonpluripolar product of closed positive $(1,1)$-current $T_{\min}$ with minimal singularities in $\alpha$, that is,
$\langle\alpha^{p}\rangle:=\{\langle T_{\min}^{p}\rangle\}\in H^{p,p}(X,\mathbb{R}).$
By [BEGZ10, Theorem 1.16], a pseudo-effective class $\alpha$ is big if $\langle\alpha^{n}\rangle>0$. Furthermore, if a big class $\alpha$ is nef, then $\langle\alpha^{p}\rangle=\alpha^{p}$ for any $p>0$ by [Mar13, Proposition 1]. The following property is related to one of our main results Corollary 1.9 (3).
Proposition 2.8 ([Bou04], Proposition 3.2, Proposition 3.6).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T=\theta+dd^{c}\varphi$ be a closed positive $(1,1)$-current with minimal singularities in $\alpha$ where $\theta$ is a smooth $(1,1)$-form in $\alpha$. Then the Lelong number of $T$ identically 0 at all point in $X$. That is,
$\inf_{\varepsilon>0}\sup_{|z-x|<\varepsilon}\frac{\varphi}{\log|z-x|}=0$
for any $x\in X$.
In this paper, we will use the following general result.
Theorem 2.9 ([BEGZ10], Theorem 3.1, Theorem 4.1, Theorem 5.1, [GPTW24]).
Let $X$ be a compact Kähler manifold and $\alpha$ be a big class on $X$. Then the following holds:
  1. $(1)$
    If $\mu$ is a nonpluripolar measure on $X$ with $\mu(X)=\langle\alpha^{n}\rangle$, then there exists a closed positive $(1,1)$-current $T$ in $\alpha$ such that $T$ satisfies the following complex Monge-Ampère equation:
    $\langle T^{n}\rangle=\mu.$ (2.1)
  2. $(2)$
    If $\mu=e^{f}dV$ where $dV$ is a smooth volume form on $X$ and $e^{f}\in L^{p}(X)$ with $p>1$, then the solution $T$ of (2.1) has minimal singularities in $\alpha$.
  3. $(3)$
    If $\alpha$ is nef and big and $f\in C^{\infty}(X)$, then the solution $T$ of (2.1) is smooth on ${\rm{Amp}}(\alpha)$.
  4. $(4)$
    Assume that $\alpha$ is a nef and big class and $\theta$ be a smooth $(1,1)$-form in $\alpha$. Let $T=\theta+dd^{c}\varphi$ be a closed positive $(1,1)$-current in $\alpha$ which satisfies $\sup_{X}\varphi=0$ and
    $\langle T^{n}\rangle=e^{f}\omega_{0}^{n}$
    where $e^{f}\in L^{p}(X)$ with $p>1$. Let $\omega_{i}:=(\theta+(1/i)\omega_{0})+dd^{c}\varphi_{i}$ be a smooth Kähler metric in a Kähler class $\alpha+(1/i)\omega_{0}$ given by the solution to
    $\omega_{i}^{n}=e^{f_{i}}\omega_{0}^{n}.$
    Assume that $f_{i}\in C^{\infty}(X)$ and $\int_{X}|f_{i}|^{p}e^{f_{i}}\omega_{0}^{n}\leq C$ for any $i$. Then, there exists a constant $C>0$ such that the following holds for any $i$:
    $0\leq-\varphi_{i}+V_{i}\leq C,$
    where $V_{i}=\sup\{\psi\mid\hbox{ nonpositive $(\theta+(1/i)\omega_{0})$-plurisubharmonic function on $X$}\}$. In particular, if $D=E_{nK}(\alpha)$ is a snc divisor, then the following holds:
    $\varphi_{i}\geq\log|s_{D}|^{2}-C.$
2.2 Definition of $\alpha^{n-1}$-slope polystability
In this section, we review the definition of slope stability for nef and big classes and its important properties. The precise definition of slope polystability for nef and big classes was firstly given by [Jin25-2, Definition 3.2]
Definition 2.10.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\mathcal{E}$ be a torsion free coherent sheaf on $X$.
  1. $(1)$
    We define the $\alpha^{n-1}$-slope of $\mathcal{E}$, denoted by $\mu_{\alpha}(\mathcal{E})$, by
    $\mu_{\alpha}(\mathcal{E}):=\frac{1}{{\rm{rk}}\mathcal{E}}\int_{X}c_{1}(\mathcal{E})\wedge\alpha^{n-1}.$
  2. $(2)$
[…]
    $\mu_{\alpha}(\mathcal{F})<\mu_{\alpha}(\mathcal{E})\hskip 5.69054pt(\hbox{ resp. }\mu_{\alpha}(\mathcal{F})\leq\mu_{\alpha}(\mathcal{E}))$
    holds.
  3. $(3)$
    A torsion free sheaf $\mathcal{E}$ is $\alpha^{n-1}$-slope polystable if there exists $\alpha^{n-1}$-slope stable torsion free sheaves $\mathcal{E}_{1},\ldots,\mathcal{E}_{k}$ on $X$ with $\mu_{\alpha}(\mathcal{E}_{1})=\cdots=\mu_{\alpha}(\mathcal{E}_{k})$ such that there exists an isomorphism
    $\mathcal{E}|_{{\rm{Amp}}(\alpha)}\simeq(\mathcal{E}_{1}\oplus\cdots\oplus\mathcal{E}_{k})|_{{\rm{Amp}}(\alpha)}$
    as coherent sheaves on ${\rm{Amp}}(\alpha)$.
In singular settings, we define $\alpha^{n-1}$-slope stability via resolution of singularities as follows.
Definition 2.11.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is a compact Kähler manifold. Then a torsion free sheaf $\mathcal{E}$ on $X$ is $\alpha^{n-1}$-slope stable (resp. semistable, polystable) if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is $(\pi^{*}\alpha)^{n-1}$-slope stable (resp. semistable, polystable).
By the following lemma, above Definition 2.11 is independent of the choice of a resolution $\pi$.
Lemma 2.12 ([IJZ25], Proposition 3.28).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\mathcal{E}$ be a torsion free sheaf on $X$. Let $\pi:Y\to X$ be a composition of blow-ups along complex submanifolds contained in $E_{nK}(\alpha)$ or of codimension at least 2. Then $\mathcal{E}$ is $\alpha^{n-1}$-slope polystable if and only if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is $(\pi^{*}\alpha)^{n-1}$-slope polystable. In particular, Definition 2.11 is independent of the choice of a resolution $\pi$.
The proof of our main Theorem 1.4 relies on the openness of $\alpha^{n-1}$-slope stability in the following sense.
Lemma 2.13 (see e.g.[Jin25-2], Lemma 4.5).
Let $X$ be a compact Kähler manifold with a Kähler class $\omega_{0}$ and $\alpha$ be a nef and big class. If a holomorphic vector bundle $E$ is $\alpha^{n-1}$-slope stable, then $E$ is $(\alpha+\varepsilon\omega_{0})^{n-1}$-slope stable for sufficiently small $\varepsilon>0$.
Remark 2.14.
In [Jin25], the author defined the notion of slope stability if $\alpha$ is big, not necessarily nef, cohomology class by replacing $\alpha^{n-1}$ by the positive product $\langle\alpha^{n-1}\rangle$ [IJZ25, Definition 3.22]. If $\alpha$ satisfies the vanishing property in the sense of [IJZ25, Definition 3.23], in particular if $X$ is Moishezon, then Lemma 2.12 also holds by [IJZ25, Proposition 3.28].
2.3 Adapted closed positive $(1,1)$-currents
In this subsection, we introduce the notion of adapted closed positive $(1,1)$-currents which plays an essential role in this paper. The following definition is inspired by the tame approximations of closed positive $(1,1)$-currents with bounded potentials by [CCHTST25].
Definition 2.15.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Then, a closed positive $(1,1)$-current $T$ in $\alpha$ is said to be adapted if it satisfies the following conditions:
  1. $(1)$
    $T$ is smooth Kähler on ${\rm{Amp}}(\alpha)$,
Let $\pi:Y\to X$ be a composition of blow-ups so that $D=E_{nK}(\pi^{*}\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$. Then
  1. $(2)$
    $\pi^{*}T\geq|s_{D}|^{2m}_{h_{D}}\omega_{Y}$ for some $m\in\mathbb{Z}_{\geq 0}$ where $\omega_{Y}$ is a smooth Kähler metric on $Y$.
  2. $(3)$
    There exists a sequence of smooth Kähler metrics $\omega_{i}$ lying in a Kähler class $\pi^{*}\alpha+(1/i)\omega_{Y}$ on $Y$ such that $\omega_{i}$ locally smoothly converges to $\pi^{*}T$ on ${\rm{Amp}}(\pi^{*}\alpha)=Y\setminus D$ and
    $\int_{Y}\left(\frac{\omega_{i}^{n}}{\omega_{Y}^{n}}\right)^{p}\omega_{Y}^{n}\leq C$
    holds for some constant $C>0$ and $p>1$ both independent of $i$. We say the above $\omega_{i}$ by an adapted approximation of $T$.
Proposition 2.16 ([Jin25-2], Lemma 4.4, see also [BEGZ10], Theorem 5.1).
Let $X$ be a compact Kähler manifold with a smooth Kähler metric $\omega_{0}$ on $X$ and $\alpha$ be a nef and big class on $X$. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ defined by $\langle T^{n}\rangle=e^{f}\omega_{0}^{n}$ with $f\in C^{\infty}(X)$. Then $T$ is an adapted current. In particular, any nef and big class on a compact Kähler manifold contains an adapted current.
By definition, the density function of the volume form of any adapted current $T$ with respect to $\omega_{0}^{n}$ lies in $L^{p}(X,\omega_{0}^{n})$ with $p>1$. In particular, the following holds:
Lemma 2.17 (refer to Theorem 2.9).
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$ such that $D=E_{nK}(\alpha)$ is a snc divisor. Let $s_{D}$ be a defining section of $D$ and $h_{D}$ be a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. Then, any adapted current $T$ in $\alpha$ has minimal singularities and it satisfies
$|s_{D}|_{h_{D}}^{k}\langle T^{n}\rangle\leq C\omega_{0}^{n}$
for some large $k>0$. Let $\theta$ be a smooth $(1,1)$-form in $\alpha$ and $\varphi\leq 0$ be a $\theta$-plurisubharmonic function on $X$ with $T=\theta+dd^{c}\varphi$. Then there exists a constant $C>0$ such that the following holds for any $i$:
$\varphi\geq\log|s_{D}|^{2}-C.$
We define adapted currents in singular settings via resolutions as follows.
Definition 2.18.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\pi:Y\to X$ be a resolution of singularities so that $Y$ is a smooth Kähler manifold with a Kähler class $\omega_{Y}$. Then, a closed positive $(1,1)$-current $T$ in $\alpha$ is adapted if $\pi^{*}T$ is an adapted closed positive $(1,1)$-current in $\pi^{*}\alpha$ on $Y$ in the sense of Definition 2.15
The following lemma shows that the above definition 2.18 is independent of the choices of resolution $\pi$. The proof is obvious.
Lemma 2.19.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\pi:Y\to X$ be a composition of blow-ups along complex submanifolds contained in $E_{nK}(\alpha)$ or of codimension at least 2. Then a closed positive $(1,1)$-current $T$ in $\alpha$ is adapted if and only if $\pi^{*}T$ in $\pi^{*}\alpha$ is adapted.
[…]
Theorem 2.20 ([CCHTST25]).
Let $X$ be a compact normal Kähler variety with log terminal singularities. Assume that $X$ admits a singular Kähler-Einstein metric $\omega$. Then $\omega$ is an adapted current in the sense of Definition 2.18.
2.4 $T$-adapted HYM metrics
Bando-Siu [BS94] introduced the notion of admissible Hermitian-Yang-Mills metrics to establish the Kobayashi-Hitchin correspondence for reflexive sheaves on compact Kähler manifolds. In this section, we introduce the notion of adapted Hermitian-Yang-Mills metrics by imposing two additional conditions on admissible Hermitian-Yang-Mills metrics.
Definition 2.21.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$. Let $\pi:Y\to X$ be a composition of blow-ups so that $D=E_{nK}(\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$. Let $(E,h_{0})$ be a complex hermitian vector bundle on $X$ with a smooth hermitian metric $h_{0}$ on $E$.
  1. $(1)$
    A $T$-admissible Hermitian-Yang-Mills connection ($T$-admissible HYM connection in short) on $(E,h_{0})$ is a smooth hermitian connection $\nabla$ on $(E,h_{0})|_{{\rm{Amp}}(\alpha)}$ which satisfies the following conditions:
    1. (a)
      (integrability) $\bar{\partial}^{\nabla}\circ\bar{\partial}^{\nabla}=0$.
    2. (b)
      (HYM equation) $\sqrt{-1}\Lambda_{T}F_{\nabla}=\lambda{\rm{id}}_{E}$ on ${\rm{Amp}}(\alpha)$ for some constant $\lambda\in\mathbb{R}$.
    3. (c)
      (admissible condition) $\int_{{\rm{Amp}}(\alpha)}|F_{\nabla}|_{h_{0},T}^{2}T^{n}<\infty$.
    We call $\lambda$ in (b) the $T$-Hermitian-Yang-Mills constant of $\nabla$.
  2. $(2)$
    If $E$ admits a holomorphic structure $\bar{\partial}^{E}$, then a $T$-admissible Hermitian-Yang-Mills metric on $(E,\bar{\partial}^{E})$ is a smooth hermitian metric $h$ on $E|_{{\rm{Amp}}(\alpha)}$ whose Chern connection $\nabla_{h}$ satisfies the following conditions:
    • (b)
      $\sqrt{-1}\Lambda_{T}F_{\nabla_{h}}=\lambda{\rm{id}}_{E}$ on ${\rm{Amp}}(\alpha)$ for some constant $\lambda\in\mathbb{R}$.
    • (c)
      $\int_{{\rm{Amp}}(\alpha)}|F_{\nabla_{h}}|_{h,T}^{2}T^{n}<\infty$.
    We call $\lambda$ in (b) the $T$-Hermitian-Yang-Mills constant of $h$.
  3. $(3)$
    If $E$ admits a holomorphic structure $\bar{\partial}^{E}$, then a $T$-adapted Hermitian-Yang-Mills metric on $(E,\bar{\partial}^{E})$ is a $T$-admissible HYM metric on $(E,\bar{\partial}^{E})$ which additionally satisfies the following two conditions: If we denote by $h=h_{0}\Psi^{2}$ with a positive definite $h_{0}$-hermitian endomorphism $\Psi$, then it satisfies the following two conditions, called by $T$-adapted conditions.
    1. (d)
      There exists $m\in\mathbb{Z}_{\geq 0}$ such that
      $\sup_{Y\setminus D}|s_{D}^{m}\pi^{*}\Psi|_{h_{D}^{m}\otimes\pi^{*}h_{0}\otimes\pi^{*}h_{0}^{*}}<\infty.$
    2. (e)
      There exists $l\in\mathbb{Z}_{\geq 0}$ such that
      $\int_{Y\setminus D}|s_{D}^{l}\partial^{\nabla_{\pi^{*}h_{0}}}(\pi^{*}\Psi)|^{2}_{h_{D}^{l}\otimes\pi^{*}h_{0}\otimes\pi^{*}h_{0}^{*},T}(\pi^{*}T)^{n}<\infty.$
We can define $T$-adapted HYM metrics in singular settings via resolution as follows. The readers can find the definition of the ample locus of big classes in singular settings from Definition 2.3.
Definition 2.22.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\mathcal{E}$ be a torsion free sheaf on $X$. Let $\pi:Y\to X$ be a resolution of singularities so that $Y$ is smooth Kähler and $E:=\pi^{*}\mathcal{E}/{\rm{Tor}}$ is locally free. Then, a $T$-adapted Hermitian-Yang-Mills metric on $\mathcal{E}$ is a smooth hermitian metric $h$ on the locally free locus of $\mathcal{E}|_{{\rm{Amp}}(\alpha)}$ such that $\pi^{*}h$ is a $\pi^{*}T$-adapted HYM metric on $E$.
By the following lemma, we can see that the Definition 2.22 is independent of the choice of a resolution $\pi$. The proof is obvious.
Lemma 2.23.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$. Let $E$ be a holomorphic vector bundle on $X$. Let $\pi:Y\to X$ be a composition of blow-ups along complex submanifolds contained in $E_{nK}(\alpha)$ or of codimension at least 2. Then a smooth hermitian metric $h$ on $E|_{{\rm{Amp}}(\alpha)}$ is a $T$-adapted HYM metric on $E$ if and only if $\pi^{*}h$ is a $\pi^{*}T$-adapted HYM metric on $\pi^{*}E$. In particular, the Definition 2.22 is independent of the choice of a resolution $\pi$.
3 Proof of the main results
In this section, we prove Theorem 1.4. We remark that, by Lemma 2.12, Lemma 2.19 and Lemma 2.23, we can assume that the non-Kähler locus of $\alpha$ is a snc divisor.
Remark 3.1.
By Lemma 2.12, Lemma 2.19 and Lemma 2.23, if one has established Theorem 1.4, then we immediately obtain its singular version Corollary 1.6.
In this section, we use the following notations.
Notation 3.2.
Let $(X,\omega_{0})$ be a compact Kähler manifold with a smooth Kähler metric $\omega_{0}$ and $\alpha$ be a nef and big class on $X$.
  • Suppose that $D=E_{nK}(\alpha)$ is an effective snc divisor. We denote by $s_{D}$ a defining section of $D$ and by $h_{D}$ a smooth hermitian metric on the holomorphic line bundle $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$.
  • Let $T$ be an adapted closed positive $(1,1)$-current in $\alpha$ satisfying $|s_{D}|^{k}\langle T^{n}\rangle\leq C\omega_{0}^{n}$ (refer to Lemma 2.17). Let $m\geq 0$ be a nonnegative integer such that $T\geq|s_{D}|^{2m}_{h_{D}}\omega_{0}$ holds in the sense of currents.
  • Let $\omega_{i}$ be a smooth Kähler metric contained in a Kähler class $\alpha+(1/i)\omega_{0}$ which gives an adapted approximation of $T$ in the sense of Definition 2.15.
  • Let $(E,h_{0})$ be a pair consisting of a holomorphic vector bundle $E$ on $X$ and a smooth hermitian metric $h_{0}$ on $E$. Assume that $E$ is $\alpha^{n-1}$-slope stable.
3.1 Preliminary results
Definition 3.3.
Let $(X,d)$ be a metric space. Then, for any subset $A\subset X$, we define the d-dimensional Hausdorff measure $\mathcal{H}^{d}(A)\in[0,\infty]$ as follows:
  1. $(1)$
    For any $\delta\in(0,\infty]$,
    $\mathcal{H}^{d}_{\delta}(A):=\inf\left\{\Sigma_{i\in I}\omega_{d}\left(\frac{{\rm diam}A_{i}}{2}\right)^{d}\mid A\subset\bigcup_{i\in I}A_{i},\hskip 5.69054pt{\rm diam}A_{i}<\delta\right\},$
    where $\omega_{d}$ is the volume of the Euclidean unit ball of dimension d and the index set $I$ is at most countable.
  2. $(2)$
    $\mathcal{H}^{d}(A):=\lim_{\delta\to 0}\mathcal{H}^{d}_{\delta}(A)$.
Furthermore, we say that $A$ has the Hausdorff dimension s if $s=\inf\{d\mid\mathcal{H}^{d}(A)<\infty\}$.
First of all, we prepare the following existence result of cut-off functions.
Lemma 3.4 ([EG15], §4.7, Theorem 3).
Let $(X,g)$ be a compact Riemannian manifold of dimension $n$ and $Z$ be a compact subset of $X$ of Hausdorff codimension 2. Then, there exists a sequence of smooth nonnegative functions $\eta_{k}:X\to[0,1]$ satisfying the following conditions:
  1. $(1)$
    ${\rm{Supp}}(\eta_{k})\Subset X\setminus Z$.
  2. $(2)$
    $\{\eta_{k}=1\}\Subset\{\eta_{k+1}=1\}$ and $\bigcup_{k=1}^{\infty}\{\eta_{k}=1\}=X\setminus Z$.
  3. $(3)$
    $\lim_{k\to\infty}\int_{X}|d\eta_{k}|_{g}^{2}{\rm{vol}}_{g}=0$.
Proof.
Although it is a direct consequence of the proof of [EG15, §4.7, Theorem 3], we include the proof for the completeness. First, we fix a finite family of coordinate covers $\{U_{i}\}_{i=1,\ldots,a}$ of $Z$. We can find a constant $C>0$ such that $C^{-1}g_{{\rm{Euc}}}<g<Cg_{{\rm{Euc}}}$ on each $U_{i}$, where $g_{{\rm{Euc}}}$ is the Euclidean metric on $U_{j}$. We fix a constant $\varepsilon_{0}>0$ and denote the $\varepsilon_{0}$-neighborhood of $Z$ by $Z_{\varepsilon_{0}}:=\{d_{g}(\cdot,Z)<\varepsilon_{0}\}$. Since $\mathcal{H}^{n-2}(Z)<\infty$ and $Z$ is compact, we can find a finite family of $g$-geodesic balls $\{B^{g}_{x_{j}}(r_{j})\}_{j=1\ldots,m}$ such that $2r_{j}<\varepsilon_{0}$, any $B^{g}_{x_{j}}(r_{j})$ is contained in some $U_{i}$, $B^{g}_{x_{j}}(r_{j})\cap Z\neq\phi$, $Z\subset\cup_{j=1}^{m}B^{g}_{x_{j}}(r_{j})$ and
$\Sigma_{j=1}^{m}r_{j}^{n-2}\leq C(\mathcal{H}^{n-2}(Z)+1)$
for some universal constant $C>0$. Now we set $W_{0}:=\bigcup_{j=1}^{m}B^{g}_{x_{j}}(r_{j})$ and define $f_{j}\in L^{\infty}(X,{\rm{vol}}_{g})$ by $f_{j}(x)=1$ if $x\in\overline{B^{g}}_{x_{j}}(r_{j})$, $f_{j}(x)=2-\frac{d_{g}(x,x_{j})}{r_{j}}$ if $x\in\overline{B^{g}}_{x_{j}}(2r_{j})\setminus B^{g}_{x_{j}}(r_{j})$, and $f_{j}(x)=0$ if $x\in X\setminus B^{g}_{x_{j}}(2r_{j})$. Since $g$ is equivalent to the Euclidean metric on each $U_{i}$, we can find a constant $C=C(g)>0$ such that $|df_{j}|\leq nCr_{j}^{-1}$. In particular, we have
$\int_{X}|df_{j}|^{2}_{g}{\rm{vol}}_{g}\leq Cn\int_{B^{g_{{\rm{Euc}}}}_{0}(r_{j})}r_{j}^{-2}{\rm{vol}}_{g_{{\rm{Euc}}}}\leq Cr_{j}^{n-2}.$
Let us define $g_{0}:=\max_{j=1,\ldots,m}f_{j}$. Then we have $W_{0}\subset\{g_{0}=1\}$, ${\rm{Supp}}(g_{0})\Subset Z_{\varepsilon_{0}}$ and
$\int_{X}|dg_{0}|_{g}^{2}{\rm{vol}}_{g}\leq\Sigma_{j=1}^{m}|df_{j}|^{2}_{g}{\rm{vol}}_{g}\leq C\Sigma_{j=1}^{m}r_{j}^{n-2}\leq C(\mathcal{H}^{n-2}(Z)+1).$
Next we define $\varepsilon_{1}:=\frac{1}{2}d_{g}(Z,X\setminus W_{0})>0$. Then, by the same way as above, we obtain $g_{1}\in L^{\infty}(X,{\rm{vol}}_{g})$ such that $Z$ is contained the internal $\{g_{1}=1\}^{o}$, ${\rm{Supp}}(g_{1})\subset Z_{\varepsilon_{1}}$ and
$\int_{X}|dg_{1}|^{2}_{g}{\rm{vol}}_{g}\leq C(\mathcal{H}^{n-2}(Z)+1).$
Inductively, we obtain a sequence of positive constants $\varepsilon_{k}\searrow 0$ and functions $g_{k}\in L^{\infty}(X,{\rm{vol}}_{g})$ such that $\overline{Z_{\varepsilon_{k+1}}}\subset\{g_{k}=1\}^{o}$, ${\rm{Supp}}(g_{k})\subset Z_{\varepsilon_{k}}$ and $\int_{X}|dg_{k}|^{2}_{g}{\rm{vol}}_{g}\leq C$. Let
$h_{j}:=\big(\Sigma_{k=1}^{j}\frac{1}{k}\big)^{-1}\Sigma_{k=1}^{j}\frac{g_{k}}{k}.$
Then $Z_{\varepsilon_{j+1}}\subset\{h_{j}=1\}^{o}$, ${\rm{Supp}}(h_{j})\subset Z_{\varepsilon_{j}}$ and
$\int_{X}|dh_{j}|^{2}_{g}{\rm{vol}}_{g}$ $=\big(\Sigma_{k=1}^{j}\frac{1}{k}\big)^{-2}\Sigma_{k=1}^{j}\frac{1}{k^{2}}\int_{X}|dg_{k}|^{2}_{g}{\rm{vol}}_{g}$
$\leq C\big(\Sigma_{k=1}^{j}\frac{1}{k}\big)^{-2}\Sigma_{k=1}^{j}\frac{1}{k^{2}}$
$\rightarrow 0\hskip 11.38109pt\hbox{in $j\to\infty$}.$
Let $F:[0,\infty)\to[0,1]$ be a smooth cut-off function such that $F=1$ on $0,\frac{1}{2}]$ and $F=0$ on $1,\infty)$. Then we can define
$\widetilde{\eta_{j}}:=F(h_{j})$
and $\eta_{j}$ a regularization of $\widetilde{\eta_{j}}$. Then $\eta_{j}$ satisfies the three conditions in the statement of this lemma. ∎
Lemma 3.5 ([Jin25-2]).
In the notations in Notation 3.2, the following holds:
  1. $(1)$
    $E$ is $\{\omega_{i}\}^{n-1}$-slope stable.
Let $h_{i}=h_{0}\Psi_{i}^{2}$ be the $\omega_{i}$-HYM metric on $E$ where $\Psi_{i}$ is a positive definite $h_{0}$-hermitian endomorphism of $E$ with $\det\Psi_{i}=1$. Then
  1. $(2)$
    $\nabla_{i}:=\Psi_{i}\circ\nabla_{h_{i}}\circ\Psi_{i}^{-1}$ is an $\omega_{i}$-HYM connection of a complex hermitian vector bundle $(E,h_{0})$.
  2. $(3)$
    $\Psi_{i}$ defines a holomorphic isomorphism from $(E,\bar{\partial}^{E})$ to $(E,\bar{\partial}^{\nabla_{i}})$.
  3. $(4)$
    There is a constant $C>0$ independent of $i$ such that the $L^{2}$-norm of the curvature tensor $F_{\nabla_{i}}$ of $\nabla_{i}$ is uniformly bounded by $C$, that is, the following holds:
    $\int_{X}|F_{\nabla_{i}}|_{h_{0},\omega_{i}}\omega_{i}^{n}\leq C.$
  4. $(5)$
    There exists an analytic subset $\Sigma$ in $X\setminus D$ of codimension at least 2 and a $T$-admissible HYM connection $\nabla_{\infty}$ on $(E,h_{0})|_{X\setminus Z}$, where $Z=\overline{D\cup\Sigma}$, such that $A_{i}:=\nabla_{i}-\nabla_{h_{0}}$, a smooth ${\rm{End}}(E)$ valued 1-form, satisfies $\|A_{i}\|_{C^{k}(A,\omega_{i},h_{0})}\to 0$ in $i\to\infty$ for any $k>0$ and any compact subset $A\subset X\setminus Z$.
In the following subsections, we construct $\Psi_{\infty}^{\pm 1}$ the nontrivial $L^{2}_{1}(X\setminus Z,T)$-local limit of $\Psi_{i}^{\pm 1}$ in Lemma 3.5 above, which gives a $T$-adapted HYM metric on $E$ (refer to Theorem 3.17).
3.2 Construction of $\Psi_{\infty}$
The following mean value type inequality by Guo-Phong-Sturm [GPS24] plays an important role.
Lemma 3.6 ([GPS24], Lemma 2).
Let $\omega_{i}$ be a smooth Kähler metric in Notation 3.2. Suppose that $v\in L^{1}(X,\omega_{i}^{n})$ is a function which satisfies $\int_{X}v\omega_{i}^{n}=0$ and
$v\in C^{2}(\overline{\Omega_{0}}),\hskip 11.38109pt\Delta_{\omega_{i}}v\geq-a\hskip 5.69054pt\hbox{on $\Omega_{0}$}$
for some constant $a>0$ and $\Omega_{s}=\{v>s\}$. Then, there exists a constant $C>0$ independent of $i$ and $v$ such that
$\sup_{X}v\leq C\big(a+\|v\|_{L^{1}(X,\omega_{i}^{n})}\big).$
3.2.1 Reduction of the construction to a uniform $C^{0}$-estimate
The following lemma shows that it suffices to derive the uniform $C^{0}$-estimates for the construction of $\Psi_{\infty}^{\pm 1}$.
Lemma 3.7.
We use the notations in Notation 3.2 and Lemma 3.5. If there exists a constant $C>0$ independent of $i$ such that
$\sup_{X}|s_{D}\Psi_{i}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C$ (3.1)
holds for any $i$, then there exists a holomorphic morphism $\Psi_{\infty}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}:=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ such that the following holds:
  1. $(1)$
    there exists a subsequence of $\{\Psi_{i}\}_{i}$, also denoted by $\{\Psi_{i}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$,
    $\lim_{i\to\infty}\int_{A}|\Psi_{i}-\Psi_{\infty}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
  2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that
    $\sup_{X}|s_{D}\Psi_{\infty}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
  3. $(3)$
    Furthermore, $\Psi_{\infty}$ is nontrivial: $\Psi_{\infty}\not\equiv 0$.
Proof.
Let $\eta_{k}:X\to[0,1]$ be a sequence of smooth functions such that ${\rm{Supp}}(\eta_{k})\Subset X\setminus Z$ and
$A_{k}:=\{\eta_{k}=1\}\nearrow X\setminus Z$ (3.2)
in $k\to\infty$. Such $\eta_{k}$ exists by Lemma 3.2.
Let us define an $h_{0}^{*}\otimes h_{0}\otimes h_{D}$-connection on ${\rm{End}}(E)\otimes\mathcal{O}(D)$ by $\widetilde{\nabla_{i}}:=\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}$. Since $\Psi_{i}:(E,\bar{\partial}^{E})\to(E,\bar{\partial}^{\nabla_{i}})$ is holomorphic, we immediately have $\bar{\partial}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})=0$. Then if we use the identity $\Box^{\partial^{\widetilde{\nabla_{i}}}}_{\omega_{i}}=\Box^{\bar{\partial}^{\widetilde{\nabla_{i}}}}_{\omega_{i}}-\sqrt{-1}[\Lambda_{\omega_{i}},F_{\widetilde{\nabla_{i}}}$], we obtain
$\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})$ $=2\Box_{\omega_{i}}^{\bar{\partial}^{\widetilde{\nabla_{i}}}}(s_{D}\Psi_{i})-\sqrt{-1}\Lambda_{\omega_{i}}F_{\widetilde{\nabla_{i}}}\cdot(s_{D}\Psi_{i})$
$=-\sqrt{-1}\Lambda_{\omega_{i}}F_{\widetilde{\nabla_{i}}}\cdot(s_{D}\Psi_{i}).$
In particular, we obtain
$\|\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})\|_{L^{2}(A_{k})}\leq C_{k}$ (3.3)
for any $i$. Next we show
$\|\widetilde{\nabla_{i}}(\eta_{k}s_{D}\Psi_{i})\|^{2}_{L^{2}(X)}\leq(C^{\prime}_{k}+1)\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}+\|\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})\|_{L^{2}(A_{2k})}^{2}\leq C_{k}.$ (3.4)
First we compute as follows:
$(\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i}),\eta_{k}^{2}s_{D}\Psi_{i})_{L^{2}}$ $=(\widetilde{\nabla_{i}}(\eta_{k}^{2}s_{D}\Psi_{i}),\widetilde{\nabla_{i}}s_{D}\Psi_{i})$
$=(2\eta_{k}\nabla\eta_{k}\cdot s_{D}\Psi_{i}+\eta_{k}^{2}\widetilde{\nabla_{i}}(s_{D}\Psi_{i}),\widetilde{\nabla_{i}}(s_{D}\Psi_{i}))$
$=2(s_{D}\Psi_{i}\nabla\eta_{k},\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i}))+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$\geq-2(2\|s_{D}\Psi_{i}\nabla\eta_{k}\|)(\frac{1}{2}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|)+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$\geq-2(4\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}+\frac{1}{4}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2})+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$=-8\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}+\frac{1}{2}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}.$
Thus, we obtain
$\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$ $\leq 2(\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i}),\eta_{k}^{2}s_{D}\Psi_{i})+16\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}$
$\leq 2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|_{L^{2}(A_{2k})}\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}+C_{k}\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}$
$\leq(1+C_{k})\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}+\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}.$
Therefore, by the assumption of this lemma and with (3.3), the following estimate follows:
$\|\widetilde{\nabla_{i}}(\eta_{k}s_{D}\Psi_{i})\|^{2}$ $\leq 2(1+C^{\prime}_{k})\|s_{D}\Psi_{i}\|^{2}_{L^{2}(A_{2k})}+2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}\leq C^{\prime\prime}_{k},$
which shows (3.4). By Lemma 3.5 (5), $\nabla_{i}$ smoothly converges to $\nabla_{\infty}$ on each $A_{k}$ and by the assumption that $\omega_{i}$ smoothly converges to $T$ on each $A_{k}$. Therefore, if we define $\widetilde{\nabla_{\infty}}:=\nabla_{\infty}\otimes\nabla_{h_{0}}\otimes\nabla_{h_{D}}$, then we obtain that
$\|\widetilde{\nabla_{\infty}}(\eta_{k}s_{D}\Psi_{i})\|^{2}_{L^{2}(X)}\leq C_{k}\|s_{D}\Psi_{i}\|^{2}_{L^{2}(A_{2k})}+2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}\leq C^{\prime}_{k},$ (3.5)
By the assumption (3.1) and above (3.5), the Sobolev embedding theorem ensures that, for any $k$, we can find a subsequence $\{i^{k}_{j}\}_{j}$ of $\{i\}_{i}$ such that $\{i^{k+1}_{j}\}_{j}\subset\{i^{k}_{j}\}_{j}$ and $\eta_{k}\Psi_{i^{k}_{j}}\xrightarrow{j\to\infty}\widetilde{\Psi_{k}}$ strongly in $L^{2}(X)$. Since $\eta_{k}=\eta_{k+1}=\cdots$ on $A_{k}=\{\eta_{k}=1\}$ by (3.2) and by the choice of the subsequence, we can assume that $\widetilde{\Psi_{k}}=\widetilde{\Psi_{k+1}}$$\cdots$ on $A_{k}$. Then we define $\Psi_{\infty}$ by $\Psi_{\infty}(x):=\widetilde{\Psi_{k}}(x)$ for $x\in A_{k}$. Then $\Psi_{\infty}$ is well-defined and it is contained in $L^{2}_{{\rm{loc}}}(X\setminus Z)$. By the diagonal argument, we can find a subsequence $\{i^{k}_{j_{k}}\}_{k}$ of $\{i\}_{i}$ such that $\Psi_{i^{k}_{j_{k}}}\xrightarrow{k\to\infty}\Psi_{\infty}$ in $L^{2}_{{\rm{loc}}}(X\setminus Z)$. Since $\Psi_{i^{k}_{j_{k}}}$ converges to $\Psi_{\infty}$ at almost everywhere points in $X\setminus Z$ , the assumption (3.1) ensures that
$\sup_{X}|s_{D}\Psi_{\infty}|_{h_{0},h_{D}}\leq C.$
Since $\bar{\partial}^{{\nabla_{i}}\otimes\nabla_{h_{0}}^{*}}\Psi_{i}=0$, we can easily see that $\bar{\partial}^{\nabla_{\infty}\otimes\nabla_{h_{0}}^{*}}\Psi_{\infty}=0$ on $X\setminus Z$ in the sense of currents. Since $\Psi_{\infty}$ is an $L^{2}$ local section on $X\setminus Z$, the standard convolution argument shows that $\Psi_{\infty}$ is holomorphic with respect to $\bar{\partial}^{{\nabla_{\infty}}\otimes\nabla_{h_{0}}^{*}}$ (e.g. [Jin25-2, proof of Proposition 4.15]). Finally we see that $\Psi_{\infty}$ is nontrivial. Since $\det\Psi_{i^{k}_{j_{k}}}\equiv 1$, we have $|\Psi_{i^{k}_{j_{k}}}|\geq 1$. Since $\Psi_{i^{k}_{j_{k}}}$ converges to $\Psi_{\infty}$ almost everywhere on $X\setminus Z$, we have $|\Psi_{\infty}|\geq 1$ almost everywhere on $X\setminus Z$. Since $\Psi_{\infty}$ is continuous on $X\setminus Z$, we have that $|\Psi_{\infty}|\geq 1$ on $X\setminus Z$. In particular, we have that $\Psi_{\infty}$ is nontrivial. ∎
3.2.2 Reduction of the uniform $C^{0}$-estimate to a uniform $L^{1}$-estimate
By the next lemma, we can reduce the uniform $C^{0}$-estimate to the uniform $L^{1}$-estimate.
Lemma 3.8.
We use the notations in Notation 3.2 and Lemma 3.5. We denote by $\Psi_{i}=e^{s_{i}}$ for some $h_{0}$-hermitian endomorphism $s_{i}$ with ${\rm{Tr}}(s_{i})=0$. Then there exists a constant $C>0$ independent of $i$ such that the following holds:
$\sup_{X}\left(\log|s_{D}|_{h_{D}}^{2}+|s_{i}|_{h_{0}}\right)\leq\sup_{X}\log|s_{D}\Psi_{i}^{\pm 1}|_{h_{0}\otimes h_{D}}^{2}\leq C(1+\int_{X}\left|\log|s_{D}|^{2}+|s_{i}|_{h_{0}}\right|\omega_{i}^{n}).$ (3.6)
Proof.
Since $\det\Psi_{i}\equiv 1$, we have $|\Psi_{i}|_{h_{0}}\geq 1$. Hence there exists a constant $a>0$ such that
$\Omega_{a}=\{\log|s_{D}\Psi_{i}|^{2}>-a\}$
is not empty for all $i$ and $\Omega_{a}$ is a relatively compact open subset in $X\setminus D$ since $|\Psi_{i}|\geq 1$ on $X$ and $|s_{D}|=0$ on $D$. Hence $\omega_{i}$ smoothly converges to $T$ on $\Omega_{a}$ and $T$ is smooth Kähler metric on $\Omega_{a}$. Then, as [Sim88, Lemma 3.1], there exists a constant $C>0$ independent of $i$ such that
$\Delta_{\omega_{i}}\left(\log|s_{D}\Psi_{i}|^{2}+a-\int_{X}\log|s_{D}\Psi_{i}|^{2}\omega_{i}^{n}\right)\geq-2|\Lambda_{\omega_{i}}F_{h_{0}}|_{h_{0}}-2|\Lambda_{\omega_{i}}F_{h_{D}}|_{h_{D}}-C_{1}\geq-C$
holds on $\Omega_{a}$. In fact, let us consider the following composition of holomorphic morphisms:
$(E,\bar{\partial}^{E})\xrightarrow{\Psi_{i}}(E,\bar{\partial}^{\nabla_{i}})\xrightarrow{s_{D}}(E\otimes\mathcal{O}(D),\bar{\partial}^{\nabla_{i}\otimes\nabla_{h_{D}}}).$
Then the following computation works on $\Omega_{a}\Subset X\setminus D$:
$\Delta_{\omega_{i}}\log|s_{D}\Psi_{i}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}}$ $=2\sqrt{-1}\Lambda_{\omega_{i}}\partial\left(|s_{D}\Psi_{i}|^{-2}\langle s_{D}\Psi_{i},\partial^{\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}}(s_{D}\Psi_{i})\rangle\right)$
$=-2|s_{D}\Psi_{i}|^{-4}\left|\langle s_{D}\Psi_{i},\partial^{\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}}(s_{D}\Psi_{i})\rangle\right|^{2}$
$\hskip 11.38109pt+2|s_{D}\Psi_{i}|^{-2}\left|\partial^{\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}}(s_{D}\Psi_{i})\right|^{2}$
$\hskip 11.38109pt-2|s_{D}\Psi_{i}|^{-2}\langle s_{D}\Psi_{i},\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}}(s_{D}\Psi_{i})\rangle$
$\geq-2|\lambda_{i}|-2|\Lambda_{\omega_{i}}F_{h_{0}}|-2|\Lambda_{\omega_{i}}F_{h_{D}}|.$
Then, by the mean value type inequality of Guo-Phong-Sturm (Lemma 3.6), we obtain that there is a positive constant $C>0$ independent of $i$ such that
$\sup_{X}\left(\log|s_{D}\Psi_{i}|^{2}+a-\int_{X}\log|s_{D}\Psi_{i}|^{2}\omega_{i}^{n}\right)$
$\leq C\left(1+\int_{X}\left|\log|s_{D}\Psi_{i}|^{2}+a-\int_{X}\log|s_{D}\Psi_{i}|^{2}\omega_{i}^{n}\right|\omega_{i}^{n}\right)$
holds. Since the constant $a>0$ is independent of $i$, we can find a constant $C>0$ independent of $i$ such that the following holds:
$\sup_{X}\left(\log|s_{D}|^{2}+\log|\Psi_{i}|^{2}\right)\leq C\big(1+\int_{X}|\log\left|s_{D}\Psi_{i}|\right|\omega_{i}^{n}\big).$ (3.7)
By applying the same argument to $\log|s_{D}\Psi^{-1}_{i}|^{2}$, we obtain
$\sup_{X}\left(\log|s_{D}|^{2}+\log|\Psi_{i}^{-1}|^{2}\right)\leq C\big(1+\int_{X}|\log\left|s_{D}\Psi_{i}^{-1}|\right|\omega_{i}^{n}\big),$ (3.8)
We denote by $\Psi_{i}=e^{s_{i}}$ for some $h_{0}$-hermitian endomorphism $s_{i}$ of $E$. Choosing a suitable $h_{0}$-orthonormal basis $e_{1},\ldots,e_{r}$ around $x\in X$ so that $s_{i}$ becomes a diagonal matrix. We denote by $s_{i}=\lambda_{1}e_{1}+\cdots+\lambda_{r}e_{r}$. Then, for any $i$, we have
$\log|\Psi_{i}^{\pm 1}|^{2}=\log(e^{\pm 2\lambda_{1}}+\cdots+e^{\pm 2\lambda_{r}})\geq\pm 2\lambda_{i}$ (3.9)
and
$\log|\Psi_{i}^{\pm 1}|^{2}$ $\leq\log(e^{2|\lambda_{1}|}+\cdots+e^{2|\lambda_{r}|})$
$\leq\log(e^{2|\lambda_{1}|+\cdots+2|\lambda_{r}|}+C)$
$\leq C(|\lambda_{1}|+\cdots+|\lambda_{r}|+1)$
$\leq C(|s_{i}|_{h_{0}}+1)$ (3.10)
Then, by (3.7), (3.8), (3.9) and (3.2.2), we obtain that
$\sup_{X}\left(\log|s_{D}|^{2}+|s_{i}|^{2}\right)\leq\sup_{X}\log|s_{D}\Psi_{i}^{\pm 1}|_{h_{0}\otimes h_{D}}^{2}\leq C(1+\int_{X}\left|\log|s_{D}|^{2}+|s_{i}|_{h_{0}}\right|\omega_{i}^{n}).$
3.2.3 Proof of the uniform $L^{1}$-estimate
Then we establish the uniform $L^{1}$-estimate.
Lemma 3.9.
We use the notations in Notation 3.2, Lemma 3.5 and Lemma 3.8. Suppose that the $L^{1}$-norm in Lemma 3.8 diverges to $\infty$, that is, we suppose that $\delta_{i}^{-1}:=\int_{X}|\log|s_{D}|+|s_{i}||\omega_{i}^{n}\nearrow\infty$ in $i\to\infty$. Then, there exists a nontrivial $h_{0}$-hermitian endomorphism $u_{\infty}$ which satisfies the following.
  1. $(1)$
    For any relatively compact open subset $K\Subset X\setminus D$, there exists a subsequence of $\{u_{i}\}_{i}$, also denoted by $\{u_{i}\}_{i}$, such that $u_{i}:=\delta_{i}s_{i}\to u_{\infty}$ weakly in $L^{2}_{1}(K,T)$ and strongly in $L^{2}(K,T^{n})$.
  2. $(2)$
    Let $\Psi:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{>0}$ be a smooth positive function such that $\Psi(a,b)<\frac{1}{b-a}$ holds for any $a<b$. Then we have
    $0\geq\int_{X\setminus D}\langle\Psi(u_{\infty})(\partial^{\nabla_{h_{0}}}u_{\infty}),\partial^{\nabla_{h_{0}}}u_{\infty}\rangle_{h_{0},T}T^{n}+\sqrt{-1}\int_{X\setminus D}{\rm{Tr}}(u_{\infty}\Lambda_{T}F_{h_{0}})T^{n}.$
    We can refer to [Sim88, Section 4] for the definition of the first term above (see also [KO25, section 2.C]).
  3. $(3)$
    The eigen values of $u_{\infty}$ are constants. Let us denote by $\lambda_{1},\ldots,\lambda_{k}$ the eigen values of $u_{\infty}$ and suppose that the multiplicity of $\lambda_{j}$ is $i_{j}$ where $i_{1}+\cdots+i_{k}=r$. Then, there exists $l\leq r$ such that $\mathcal{F}_{k}:={\rm{Ker}}(u_{\infty}-\Sigma_{j=1}^{l}\lambda_{j}I_{i_{j}})$ produces a coherent subsheaf of $E$ on $X$ and it satisfies
    $\mu_{\alpha}(\mathcal{F}_{k})\geq\mu_{\alpha}(E).$
In particular, there exists a uniform constant $C>0$ such that the following holds:
$\int_{X}\left(-\log|s_{D}|^{2}+|s_{i}|\right)\omega_{i}^{n}\leq C.$
Proof.
The proof of this lemma is inspired by the argument of [KO25, Claim 4.6]. We will divide the proof into four-steps as [KO25, Claim 4.6].
1st Step: This is the only step that requires an argument different from that in [KO25, Claim 4.6]. Let $\Phi(x,y):=\frac{\exp(x-y)-1}{x-y}$ for $x,y\in\mathbb{R}$. Then, since $h_{i}=h_{0}e^{s_{i}}$ defines an $\omega_{i}$-HYM metric on $E$ and $u_{i}=\delta_{i}s_{i}$, we have (refer to [KO25, Lemma 2.9])
$0=\delta_{i}^{-1}\int_{X}\langle\Phi(u_{i}/\delta_{i})(\partial^{\nabla_{h_{0}}}u_{i}),\partial^{\nabla_{h_{0}}}u_{i}\rangle_{h_{0},\omega_{i}}\omega_{i}^{n}+\int_{X}{\rm{Tr}}(u_{i}\sqrt{-1}\Lambda_{i}F_{h_{0}})\omega_{i}^{n}.$ (3.11)
We then prove $\int_{X}{\rm{Tr}}(u_{i}\sqrt{-1}\Lambda_{\omega_{i}}F_{h_{0}})\omega_{i}^{n}$ is uniformly bounded from above and below. Since $h_{0}$ is a smooth hermitian metric on $E$ and $\omega_{0}$ is a smooth Kähler metric on $X$, there is a constant $C>0$ such that
$C\omega_{0}\cdot{\rm{Id}}_{E}\leq\sqrt{-1}F_{h_{0}}\leq C\omega_{0}\cdot{\rm{Id}}_{E}.$
By Lemma 3.8 and the definition of $u_{i}$, we have
$|u_{i}|_{h_{0}}\leq C-C\delta_{i}\log|s_{D}|_{h_{D}}^{2}.$ (3.12)
Then we have
$|u_{i}|_{h_{0}}|\sqrt{-1}\Lambda_{\omega_{i}}F_{h_{0}}|_{h_{0}}\omega_{i}^{n}\leq C(C-\delta_{i}\log|s_{D}|^{2})\omega_{0}\wedge\omega_{i}^{n-1}.$ (3.13)
By Claim 3.10 below, we obtain
$\left|\delta_{i}^{-1}\int_{X}\langle\Phi(u_{i}/\delta_{i})(\partial^{\nabla_{h_{0}}}u_{i}),\partial^{\nabla_{h_{0}}}u_{i}\rangle_{h_{0},\omega_{i}}\omega_{i}^{n}\right|\leq\int_{X}|u_{i}|_{h_{0}}|\sqrt{-1}\Lambda_{\omega_{i}}F_{h_{0}}|_{h_{0}}\omega_{i}^{n}\leq C.$ (3.14)
Claim 3.10.
There exists a constant $C>0$ such that the following holds for any $i$:
$\int_{X}(-\log|s_{D}|_{h_{D}}^{2})\wedge\omega_{0}\wedge\omega_{i}^{n-1}\leq C.$
proof of claim.
Let $\theta$ be a smooth $(1,1)$-form in $\alpha$ and $\varphi_{i}\leq 0$ be a $(\theta+(1/i)\omega_{0})$-plurisubharmonic function such that $\omega_{i}=\theta_{i}+dd^{c}\varphi_{i}$ where $\theta_{i}=\theta+(1/i)\omega_{0}$. We denote by $\omega_{D}$ the curvature form of $h_{D}$. Then
$\int_{X}(-\log|s_{D}|^{2})\wedge\omega_{0}\wedge\omega_{i}^{n-1}$
$=\int_{X}(-\log|s_{D}|^{2})\wedge\omega_{0}\wedge\theta_{i}^{n-1}+\Sigma_{p=1}^{n-1}c_{p}\int_{X}(-\log|s_{D}|^{2})\wedge\omega_{0}\wedge\theta_{i}^{n-1-p}\wedge(dd^{c}\varphi_{i})^{p}$ (3.15)
where $c_{p}>0$ is a constant depending only on $p$. Since $\theta_{i}^{n-1}\leq C\omega_{0}^{n-1}$, we have
$\int_{X}(-\log|s_{D}|^{2})\wedge\omega_{0}\wedge\theta_{i}^{n-1}\leq C\int_{X}(-\log|s_{D}|^{2})\wedge\omega_{0}^{n}\leq C.$ (3.16)
We then prove the boundedness of the second term in RHS of (3.2.3) by induction on $p$. Assume $p=1$. Let $a>0$ be any positive constant. Then
$\int_{X}(-\max{\{\log|s_{D}|^{2},-a\}})\wedge\omega_{0}\wedge\theta_{i}^{n-2}\wedge dd^{c}\varphi_{i}$
$=\int_{X}\varphi_{i}(-dd^{c}\max{\{\log|s_{D}|^{2},-a\}})\wedge\omega_{0}\wedge\theta_{i}^{n-2}$
$=\int_{X}\varphi_{i}\cdot\omega_{D}\wedge\omega_{0}\wedge\theta_{i}^{n-2}$
$\leq C\int_{X}(-\log|s_{D}|^{2})\omega_{0}^{n}$
$\leq C$ (3.17)
In the third line, we used $\varphi_{i}\omega_{D}\wedge\theta_{i}^{n-2}\leq-C\log|s_{D}|^{2}\omega_{0}^{n-1}$, which is proved by $\log|s_{D}|^{2}-C\leq\varphi_{i}\leq 0$ (see Lemma 2.17) and $C\omega_{0}^{n-1}\leq\omega_{D}\wedge\theta_{i}^{n-2}\leq C\omega_{0}^{n-1}.$ Then the induction on $p$ works. We only prove the case $p=2$.
$\int_{X}(-\max{\{\log|s_{D}|^{2},-a\}})\wedge\omega_{0}\wedge\theta_{i}^{n-3}\wedge(dd^{c}\varphi_{i})^{2}$
$=\int_{X}\varphi_{i}(-dd^{c}\max{\{\log|s_{D}|^{2},-a\}})\wedge\omega_{0}\wedge\theta_{i}^{n-3}\wedge(dd^{c}\varphi_{i})$
$=\int_{(\log|s_{D}|^{2}>-a)}\varphi_{i}\cdot\omega_{D}\wedge\omega_{0}\wedge\theta_{i}^{n-3}\wedge(\theta_{i}+dd^{c}\varphi_{i})-\int_{(\log|s_{D}|^{2}>-a)}\varphi_{i}\cdot\omega_{D}\wedge\omega_{0}\wedge\theta_{i}^{n-2}$
$\leq C\int_{X}(-\max{\{\log|s_{D}|^{2},-a\}})\omega_{0}^{n-1}\wedge(\theta_{i}+dd^{c}\varphi_{i})+C\int_{X}(-\log|s_{D}|^{2})\omega_{0}^{n}$
$\leq C$ (3.18)
In the third line, we used $\varphi_{i}\omega_{D}\wedge\theta_{i}^{n-2}\leq-C\log|s_{D}|^{2}\omega_{0}^{n-1}$ as before and $\theta_{i}+dd^{c}\varphi_{i}\geq 0$. And the fourth line is the result of the case $p=1$. We end the proof of Claim 3.10. ∎
Since the rest of the proof of (1) and (2) is the same with [KO25, Claim 4.6, Step 2, Step 3, Step 4], we only describe the outline.
2nd Step: In this step, we show that $u_{i}$ subsequencially weakly converges to $u_{\infty}$ in $L^{2}_{1,{\rm{loc}}}(X\setminus D,T)$-topology. By (3.12), the eigenvalues of $u_{i}$ on a relatively compact open subset $K\Subset X\setminus D$ is contained in an interval $\alpha,\beta]$ for any $i$. We recall that $\Phi(x,y)=\frac{{\rm{exc}}(x-y)-1}{x-y}$ defined in the 1st Step satisfies $\Phi(x,y)\geq\frac{{\rm{exc}}(\alpha-\beta)-1}{\alpha-\beta}$ for any $x,y\in[\alpha,\beta]$ (refer to [KO25, Lemma 2.7]). Then we have
$\langle\Phi(u_{i})\partial^{\nabla_{h_{0}}}u_{i},\partial^{\nabla_{h_{0}}}u_{i}\rangle_{h_{0},\omega_{i}}\geq C_{K}^{-1}|\partial^{\nabla_{h_{0}}}u_{i}|_{h_{0},\omega_{i}}^{2}\hbox{ on $K$}.$
By $\delta_{i}^{-1}=\int_{X}\left|\log|s_{D}|^{2}+|s_{i}|\right|\omega_{i}^{n}\geq 1$ and (3.14), we obtain
$\int_{K}|\partial^{\nabla_{h_{0}}}u_{i}|_{h_{0},\omega_{i}}^{2}\omega_{i}^{n}\leq C_{K}\delta_{i}^{-1}\int_{X}\langle\Phi(\delta_{i}^{-1}u_{i})\partial^{\nabla_{h_{0}}}u_{i},\partial^{\nabla_{h_{0}}}u_{i}\rangle_{h_{0},\omega_{i}}\omega_{i}^{n}\leq C_{K}.$ (3.19)
Let us consider a sequence of relatively compact open subsets in $X\setminus D$ by $X\setminus D_{\varepsilon_{i}}$ where $D_{\varepsilon_{i}}$ is an $\varepsilon_{i}$-neighborhood of $D$ with respect to $\omega_{0}$ with $\varepsilon_{i}\searrow 0$. We recall that $\omega_{i}$ smoothly converges to $T$ on each $X\setminus D_{\varepsilon_{i}}$ and $T$ is a smooth Kähler metric on $X\setminus D$. Then by (3.12) and (3.19), the standard diagonal procedure gives a subsequence of $\{u_{i}\}_{i}$, also denoted by $\{u_{i}\}_{i}$, such that
$u_{i}\to u_{\infty}\hbox{ weakly in $L^{2}_{1,{\rm{loc}}}(X\setminus D,T)$ and strongly in $L^{2}_{{\rm{loc}}}(X\setminus D,T^{n})$}.$ (3.20)
By (3.13) and Claim 3.10, the dominant convergence theorem gives
$\lim_{i\to\infty}\sqrt{-1}\int_{X}{\rm{Tr}}(u_{i}\Lambda_{\omega_{i}}F_{h_{0}})\omega_{i}^{n}\ =\sqrt{-1}\int_{X\setminus D}{\rm{Tr}}(u_{\infty}\Lambda_{T}F_{h_{0}})T^{n}.$ (3.21)
3rd Step: In this step, we show that $u_{\infty}$, the weak limit of $u_{i}$ constructed in 2nd step, is not identically 0. Since $\delta_{i}=\left(\int_{X}(-\log|s_{D}|^{2}+|s_{i}|)\omega_{i}^{n}\right)^{-1}$ and $u_{i}=\delta_{i}s_{i}$, we have
$\int_{X}\delta_{i}\log|s_{D}|^{2}\omega_{i}^{n}+\int_{X}|u_{i}|\omega_{i}^{n}=1.$
If we recall that $\int_{X}(-\log|s_{D}|)\omega_{i}^{n}\leq C$, then we obtain
$2\int_{X}(-\delta_{i}\log|s_{D}|^{2})\omega_{i}^{n}+\int_{X}|u_{i}|\omega_{i}^{n}\geq 1-\delta_{i}C.$ (3.22)
By (3.12) and again by $\int_{X}(-\log|s_{D}|)\omega_{i}^{n}\leq C$, if we take the limit as $i\to\infty$, the dominant convergence theorem allows us to interchange the limit and the integral in the LHS of (3.22). Then we obtain
$\int_{X\setminus D}|u_{\infty}|T^{n}\geq 1.$
Hence, if $u_{\infty}=0$, then it implies $0\geq 1$ which is a contradiction. Thus $u_{\infty}\neq 0$.
4th Step: In this step, we show the inequality in (2) in the statement. Let $\Psi:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{>0}$ be a smooth positive function such that $\Psi(a,b)<\frac{1}{b-a}$ when $a<b$. We fix $K\subset X\setminus D$ a relatively compact open subset. We recall that $\delta_{i}^{-1}\Phi(\delta_{i}^{-1}a,\delta_{i}^{-1}b)$ tends to $\frac{1}{b-a}$ when $a<b$ and to $\infty$ when $b\leq a$ (refer to [KO25, Lemma 2.7]). We also recall that, by (3.12), the eigenvalues of $u_{i}$ lies in an interval $\alpha,\beta]$ for all $i$. Hence, for sufficiently large $i>0$, we have
$\delta_{i}^{-1}\Phi(\delta_{i}^{-1}a,\delta_{i}^{-1}b)\geq\Psi(a,b)\hbox{ for all $a,b\in[\alpha,\beta]$}.$
Together with (3.11), we obtain
$0\geq\sqrt{-1}\int_{X}{\rm{Tr}}(u_{i}\Lambda_{\omega_{i}}F_{h_{0}})\omega_{i}^{n}+\int_{K}\langle\Psi(u_{i})\partial^{\nabla_{h_{0}}}u_{i},\partial^{\nabla_{h_{0}}}u_{i}\rangle_{h_{0},\omega_{i}}\omega_{i}^{n}.$ (3.23)
By (3.12) and (3.20), we can see that $\Psi^{\frac{1}{2}}(u_{i})\to\Psi^{\frac{1}{2}}(u_{\infty})$ in the operator norm topology on ${\rm{Hom}}(L^{2}(K,T^{n}),L^{q}(K,T^{n}))$ for any $1\leq q<2$ ([Sim88, Proposition 4.1], see also [KO25, Lemma 2.6]). Thus, by (3.21) and (3.23), we obtain that for any $\delta>0$, there exists $i_{0}>0$ such that for any $i>i_{0}$ the following holds:
$2\delta\geq\|\Psi^{\frac{1}{2}}(u_{\infty})\partial^{\nabla_{h_{0}}}\|_{L^{q}(K,T)}^{2}+\sqrt{-1}\int_{X\setminus D}{\rm{Tr}}(u_{\infty}\Lambda_{T}F_{h_{0}})T^{n}$ (3.24)
Since $\partial^{\nabla_{h_{0}}}u_{i}\to\partial^{\nabla_{h_{0}}}u_{\infty}$ weakly on $K$ (refer to (3.20)), we can again use [Sim88, Proposition 4.1] (see also [KO25, Lemma 2.6]) to obtain that $\Psi(u_{\infty})\partial^{\nabla_{h_{0}}}u_{i}\to\Psi(u_{\infty})\partial^{\nabla_{h_{0}}}u_{\infty}$ weakly in $L^{q}(K,T)$. Then, by (3.24), we obtain
$2\delta\geq\|\Psi^{\frac{1}{2}}(u_{\infty})\partial^{\nabla_{h_{0}}}u_{\infty}\|_{L^{q}(K,T)}^{2}+\sqrt{-1}\int_{X\setminus D}{\rm{Tr}}(u_{\infty}\Lambda_{T}F_{h_{0}})T^{n}.$ (3.25)
Here we recall that $1\leq q<2$, $K\Subset X\setminus D$ and $\delta>0$ are chosen arbitrary. Thus (3.25) applies when $q=2$, $K=X\setminus D$ and $\delta=0$, that is, we obtain
$0\geq\int_{X\setminus D}\langle\Psi(u_{\infty})\partial^{\nabla_{h_{0}}}u_{\infty},\partial^{\nabla_{h_{0}}}u_{\infty}\rangle_{h_{0},T}+\sqrt{-1}\int_{X\setminus D}{\rm{Tr}}(u_{\infty}\Lambda_{T}F_{h_{0}})T^{n}.$
(3) By $(1)$ and $(2)$, using Proposition 3.12 below, we can prove $(3)$ by the same argument with [Sim88, Section 5] (refer to the comment below [FO25, Claim 4.6]). We only sketch the proof. By [Sim88, Lemma 5.5], we see that the eigenvalues $\lambda_{1},\ldots,\lambda_{r}$ of $u_{\infty}$ are all constants and $\lambda_{i}$ are not all equal. Let $\{\gamma\}$ be the set of intervals between the eigenvalues of $u_{\infty}$ and for each $\gamma$, we choose $p_{\gamma}:\mathbb{R}\to\mathbb{R}$ with $p_{\gamma}(\lambda_{i})=1$ for $\lambda_{i}<\gamma$ and $p_{\gamma}(\lambda_{i})=0$ for $\lambda_{i}>\gamma$. Define $\pi_{\gamma}:=p_{\gamma}(u_{\infty})$. Then, by [Sim88, Lemma 5.6], we can see that $\pi_{\gamma}$ defines a $T$-weak holomorphic subbundle (see Definition 3.11 below). By [Sim88, Lemma 5.7], there exists $\gamma$ such that
$\mu_{\alpha}(E)\leq\frac{1}{{\rm{Tr}}(\pi_{\gamma})}\int_{X\setminus W}\left({\rm{Tr}}(\pi_{\gamma}\sqrt{-1}F_{h_{0}})-\sqrt{-1}{\rm{Tr}}(\bar{\partial}\pi_{\gamma}\wedge(\bar{\partial}\pi_{\gamma})^{*_{h_{0}}})\right)\wedge T^{n-1}$
where $W\subset X$ is an analytic subset in $X$. By Proposition 3.12 below, there exists a torsion free subsheaf $\mathcal{F}\subset E$ on $X$ such that
$\frac{1}{{\rm{Tr}}(\pi_{\gamma})}\int_{X\setminus W}\left({\rm{Tr}}(\pi_{\gamma}\sqrt{-1}F_{h_{0}})-\sqrt{-1}{\rm{Tr}}(\bar{\partial}\pi_{\gamma}\wedge(\bar{\partial}\pi_{\gamma})^{*_{h_{0}}})\right)\wedge T^{n-1}=\mu_{\alpha}(\mathcal{F}).$
It contradicts to the assumption that $E$ is $\alpha^{n-1}$-slope stable. Hence we obtain a uniform estimate $C>0$ such that the following holds:
$\int_{X}\left(-\log|s_{D}|^{2}+|s_{i}|\right)\omega_{i}^{n}\leq C.$
Definition 3.11.
Let $X$ be a compact Kähler manifold and $T$ be a closed positive $(1,1)$-current on $X$ which is smooth Kähler on $X\setminus D$ where $D$ is a complex analytic subset in $X$. Let $(E,h_{0})$ be a holomorphic hermitian vector bundle on $X$. Then, a $T$-weak holomorphic subbundle of $E$ is an $h_{0}$-hermitian endomorphism $\pi$ defined almost everywhere on $X\setminus D$ such that $\pi^{2}=\pi$, $(1-\pi)\circ\bar{\partial}^{E}\pi=0$ and
$\int_{X\setminus D}|\pi|_{h_{0}}^{2}T^{n}+\int_{X\setminus D}|\nabla_{h_{0}}\pi|^{2}_{h_{0},T}T^{n}<\infty.$
By [UY86], the image of a $T$-weak holomorphic subbundle $\pi$ uniquely defines a coherent subsheaf $\mathcal{F}$ of $E|_{X\setminus D}$, We denote $\mathcal{F}$ by ${\rm im}(\pi)$.
Following proposition is the Chern-Weil formula for $T$-weak holomorphic subbundles.
Proposition 3.12.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$ with $D=E_{nK}(\alpha)$ is a snc divisor. Let $T$ be a closed positive $(1,1)$-current with minimal singularities in $\alpha$ which is smooth Kähler on $X\setminus D$. Let $(E,h_{0})$ be a holomorphic hermitian vector bundle on $X$. Let $\pi$ be a $T$-weak holomorphic subbundle of $E$. Then the following holds.
  1. $(1)$
    There exists a torsion free coherent subsheaf $\mathcal{F}\subset E$ on $X$ such that $\mathcal{F}|_{X\setminus D}={\rm im}(\pi)$.
  2. $(2)$
    There exists an analytic subset $W\subset X$ such that
    $\int_{X}c_{1}(\mathcal{F})\wedge\alpha^{n-1}=\int_{X\setminus W}\left({\rm{Tr}}(\pi\sqrt{-1}F_{h_{0}})-{\rm{Tr}}(\bar{\partial}\pi\wedge(\bar{\partial}\pi)^{*_{h_{0}}})\right)\wedge\langle T^{n-1}\rangle.$
Proof.
Let $\pi$ be a $T$-weak holomorphic subbundle of $(E,h_{0})$ and ${\rm im}(\pi)$ be the torsion free coherent sheaf of $E|_{X\setminus D}$ defined by the image of $\pi$. Let us consider the following short exact sequence of torsion free sheaves on $X\setminus D$:
$0\to\mathcal{F}\xrightarrow{\iota}E|_{X\setminus D}\xrightarrow{p}\mathcal{G}\to 0.$ (3.26)
Then $p$ defines an $\mathcal{O}_{X}$-linear morphism $\widetilde{p}:=i_{*}p:E\to i_{*}\mathcal{G}$ where $i:X\setminus D\hookrightarrow X$ is a natural inclusion. Then $\widetilde{\mathcal{G}}:={\rm im}(\widetilde{p})/{\rm{Tor}}$ is a torsion free coherent sheaf on $X$. In fact, the dual of the surjection $\widetilde{p}:E\to{\rm im}(\widetilde{p})/{\rm{Tor}}$ gives an injection of $\mathcal{O}_{X}$-module $\widetilde{p}^{*}:({\rm im}(\widetilde{p})/{\rm{Tor}})^{*}\to E^{*}$. Then, by Oka’s coherence theorem, we obtain that $({\rm im}(\widetilde{p})/{\rm{Tor}})^{*}$ is a coherent sheaf on $X$. Hence ${\rm im}(\widetilde{p})/{\rm{Tor}}$ extends to a torsion free coherent subsheaf $\widetilde{\mathcal{G}}$ of $({\rm im}(\widetilde{p})/{\rm{Tor}})^{**}$ on $X$. Then we obtain the following short exact sequence of torsion free coherent sheaves on $X$:
$0\to\widetilde{\mathcal{F}}\xrightarrow{\widetilde{\iota}}E\xrightarrow{\widetilde{p}}\widetilde{\mathcal{G}}\to 0$ (3.27)
where the restriction of (3.27) to $X\setminus D$ coincides with (3.26). We can find an analytic subset $W\subset X$ such that ${\rm im}(\widetilde{\iota})|_{X\setminus W}\subset E|_{X\setminus W}$ is a holomorphic subbundle. Let $\widetilde{\pi}$ be an $h_{0}$-orthogonal projection from $E|_{X\setminus W}$ to $\widetilde{\mathcal{F}}|_{X\setminus W}$. Then we have
$\widetilde{\pi}|_{X\setminus W}=\pi\hbox{ almost everywhere in $X\setminus D$}.$ (3.28)
Hence, it suffices to show the following:
$\int_{X}c_{1}(\widetilde{\mathcal{F}})\wedge\alpha^{n-1}=\int_{X\setminus W}c_{1}(\widetilde{\mathcal{F}},h_{0}|_{\widetilde{\mathcal{F}}})\wedge\langle T^{n-1}\rangle,$ (3.29)
since
$c_{1}(\widetilde{\mathcal{F}},h_{0}|_{\widetilde{\mathcal{F}}})={\rm{Tr}}(\widetilde{\pi}\sqrt{-1}F_{h_{0}})-{\rm{Tr}}(\bar{\partial}\widetilde{\pi}\wedge(\bar{\partial}\widetilde{\pi})^{*_{h_{0}}}).$ (3.30)
Since (3.29) is proved by the standard way (e.g. [Kob87, Remark 8.5]), we only describe the outline of the proof. We denote by $q:={\rm{rk}}\widetilde{\mathcal{F}}$. Let us consider
$\widetilde{j}:=\wedge^{q}\widetilde{\iota}:\det\widetilde{\mathcal{F}}\to\bigwedge^{q}E.$
Then $\{\widetilde{j}=0\}=W$. Since $\widetilde{j}$ is an injection of a holomorphic line subbundle, the pull back
$u:=\widetilde{j}^{*}(\wedge^{q}h_{0})$
defines a smooth hermitian metric on $\det\widetilde{\mathcal{F}}|_{X\setminus W}$. Let $\tau$ be a local holomorphic frame of $\det\widetilde{\mathcal{F}}$ and $s_{1},\ldots,s_{r}$ be a local holomorphic frame of $E$. Then we can denote by
$\widetilde{j}(\tau)=\Sigma_{I}\tau^{I}s_{I}\hbox{ where $s_{I}=s_{i_{1}}\wedge\ldots\wedge s_{i_{p}}$ with $i_{1}<\cdots<i_{p}$}.$
Let $\widetilde{u}$ be a smooth hermitian metric on $\det\widetilde{\mathcal{F}}$ which is globally defined over $X$. Then we define a nonnegative function
$f:=\frac{u(\tau,\tau)}{\widetilde{u}(\tau,\tau)}=\Sigma_{I,J}u_{IJ}\tau^{I}\overline{\tau^{J}}$
where $u_{IJ}=\wedge^{q}h_{0}(s_{I},s_{J})/\widetilde{u}(\tau,\tau)$. Then we can see that $f$ is independent of the choice of $\tau$, that is, $f$ is globally defined on $X$, and $\{f=0\}=W$. Let $\mu:Y\to X$ be a composition of blow-ups such that $\mu^{-1}(W)$ is a snc divisor in $Y$. Since $\mu^{*}f$ vanishes along $\mu^{-1}(W)$, we have
$dd^{c}\log\mu^{*}f=[W^{\prime}]$
where $W^{\prime}]$ is an integral current supported on $\mu^{-1}(W)$ with nonnegative coefficients. We obtain that
$\int_{X\setminus(D\cup W)}dd^{c}\log f\wedge T^{n-1}=\int_{Y\setminus\mu^{-1}(D)\cup\mu^{-1}(W)}dd^{c}\log\mu^{*}f\wedge(\mu^{*}T)^{n-1}=0.$ (3.31)
Since $c_{1}(\det\widetilde{\mathcal{F}},\widetilde{u})=c_{1}(\det\widetilde{\mathcal{F}},u)+dd^{c}\log f$ on $X\setminus W$, we obtain
$\int_{X}c_{1}(\det\widetilde{\mathcal{F}})\wedge\alpha^{n-1}$ $=\int_{X}c_{1}(\det\widetilde{\mathcal{F}},\widetilde{u})\wedge\langle T^{n-1}\rangle$
$=\int_{X\setminus(D\cup W)}c_{1}(\det\widetilde{\mathcal{F}},\widetilde{u})\wedge T^{n-1}$
$=\int_{X\setminus(D\cup W)}c_{1}(\det\widetilde{\mathcal{F}},u)\wedge T^{n-1}$
$=\int_{X\setminus W}c_{1}(\det\widetilde{\mathcal{F}},u)\wedge\langle T^{n-1}\rangle.$ (3.32)
In the third line above, we used (3.31). Since $u=\widetilde{j}^{*}(\wedge^{q}h_{0})$, we have
$c_{1}(\det\widetilde{\mathcal{F}},u)=c_{1}(\widetilde{\mathcal{F}},h)\hbox{ on $X\setminus W$}.$ (3.33)
Then by (3.28), (3.30), (3.2.3) and (3.33), we obtain
$\int_{X}c_{1}(\widetilde{\mathcal{F}})\wedge\alpha^{n-1}$ $=\int_{X\setminus W}c_{1}(\widetilde{\mathcal{F}},h_{0}|_{\widetilde{\mathcal{F}}})\wedge\langle T^{n-1}\rangle$
$=\int_{X\setminus(D\cup W)}\left({\rm{Tr}}(\pi\sqrt{-1}F_{h_{0}})-{\rm{Tr}}(\bar{\partial}\pi\wedge(\bar{\partial}\pi)^{*_{h_{0}}})\right)\wedge T^{n-1}.$
3.2.4 Construction of $\Psi_{\infty}$
As a consequence of the previous subsections, we obtain the followings
Theorem 3.13.
We use the notations Notation 3.2 and Lemma 3.5. Then there exists a holomorphic morphism ${\Psi_{\infty}}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ which satisfies the following conditions:
  1. $(1)$
    There exists a subsequence of $\{{\Psi_{i}}\}_{i}$, also denoted by $\{{\Psi_{i}}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$, the following holds:
    $\lim_{i\to\infty}\int_{A}|{\Psi_{i}}-{\Psi_{\infty}}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
  2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that the following holds for any $i$:
    $\sup_{X}|s_{D}{\Psi_{i}}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
  3. $(3)$
    Furthermore, $\Psi_{\infty}$} is nontrivial: ${\Psi_{\infty}}\not\equiv 0$.
Proof.
Let us denote by $\Psi_{i}:=e^{s_{i}}$ for some $h_{0}$-hermitian endomorphism of $(E,h_{0})$ with ${\rm{Tr}}(s_{i})=0$. Then, by Lemma 3.9, we obtain that
$\int_{X}\left(-\log|s_{D}|^{2}+|s_{i}|_{h_{0}}\right)\leq C.$
Then, by Lemma 3.8, we have
$\sup_{X}|s_{D}\Psi_{i}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
Hence, by Lemma 3.7, we obtain $\Psi_{\infty}$ as in the statement. ∎
By Lemma 3.8 and Lemma 3.9, we can see that the same argument with Theorem 3.13 applies to $\Psi_{i}^{-1}$, and thus we obtain the following theorem:
Theorem 3.14.
We use the notations Notation 3.2 and Lemma 3.5. Then there exists a holomorphic morphism $\Phi_{\infty}:E_{\infty}=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ which satisfies the following conditions:
  1. $(1)$
    There exists a subsequence of $\{\Psi_{i}^{-1}\}_{i}$, also denoted by $\{\Psi_{i}^{-1}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$, the following holds:
    $\lim_{i\to\infty}\int_{A}|\Psi_{i}^{-1}-\Phi_{\infty}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
  2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that the following holds for any $i$:
    $\sup_{X}|s_{D}\Psi_{i}^{-1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
  3. $(3)$
    Furthermore, $\Phi_{\infty}$ is nontrivial: $\Phi_{\infty}\not\equiv 0$.
As a corollary of Theorem 3.13 and Theorem 3.14, we can derive the following uniform estimate on the first derivatives.
Proposition 3.15.
We use notations in Notation 3.2. Let $\Psi_{i}$}, $\Psi_{\infty}$}, $\Phi_{\infty}$ be holomorphic morphisms constructed in Theorem 3.13 and Theorem 3.14. There exists a constant $C>0$ such that the following inequality holds for any $i$:
$\int_{X}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}{\Psi_{i}}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},\omega_{i}}\omega_{i}^{n}\leq C.$
The same estimate holds for $\Psi_{i}^{-1}$. In particular, we have
$\int_{X\setminus Z}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}{\Psi_{\infty}}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},T}T^{n}<\infty.$
The same estimate holds for $\Phi_{\infty}$.
Proof.
Let $h_{D}$ be a smooth hermitian metric on $\mathcal{O}(D)$. We denote by $\nabla_{D}:=\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}$ a $h_{0}^{*}\otimes h_{0}\otimes h_{D}$-connection on ${\rm{Hom}}(E,E)\otimes\mathcal{O}(D)$. If we use the identity $\Box^{\partial^{\nabla_{D}}}_{\omega_{i}}=\Box^{\bar{\partial}^{\nabla_{D}}}_{\omega_{i}}-\sqrt{-1}[\Lambda_{\omega_{i}},F_{\nabla_{D}}]$, and $T\geq|s_{D}|^{2m}\omega_{0}$, the following calculation works:
$\int_{X}\langle\nabla_{D}(s_{D}^{m+1}{\Psi_{i}}),\nabla_{D}(s_{D}^{m+1}{\Psi_{i}})\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D},\omega_{i}}\omega_{i}^{n}$
$=\int_{X}\langle\Delta^{\nabla_{D}}_{\omega_{i}}(s_{D}^{m+1}{\Psi_{i}}),s_{D}^{m+1}{\Psi_{i}}\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}\omega_{i}^{n}$
$=-\int_{X}\langle\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{D}}\cdot(s_{D}^{m+1}{\Psi_{i}}),s_{D}^{m+1}{\Psi_{i}}\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}\omega_{i}^{n}$
$=\int_{X}\langle-\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{h_{0}}\otimes\nabla_{h_{D}}}\circ(|s_{D}|^{2m}s_{D}{\Psi_{i}}),s_{D}{\Psi}\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}\omega_{i}^{n}$
$\hskip 11.38109pt+\int_{X}\langle(|s_{D}|^{2m}s_{D}{\Psi_{i}})\circ(\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}),s_{D}{\Psi_{i}}\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}\omega_{i}^{n}$
$\leq C\int_{X}\langle|s_{D}|^{2m}\sqrt{-1}\Lambda_{\omega_{i}}\omega_{0}\cdot(s_{D}{\Psi_{i}}),s_{D}{\Psi_{i}}\rangle_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}\omega_{i}^{n}$
$\hskip 11.38109pt+|\lambda_{i}|\int_{X}|s_{D}{\Psi_{i}}|_{h_{0}^{*}\otimes h_{0}\otimes h_{D}}^{2}\omega_{i}^{n}$
$\leq C.$ (3.34)
Since $\sup_{X\setminus D}|s_{D}\Psi_{i}|\leq C$ by Theorem 3.13, we obtain that
$\int_{X}|s_{D}^{m+1}(\nabla_{i}\otimes\nabla_{h_{0}}^{*})\Psi_{i}|^{2}\omega_{i}^{n}\leq C.$
Since $\nabla_{i}=\Psi_{i}\circ\nabla_{h_{i}}\circ\Psi_{i}^{-1}$ by Lemma 3.5, the result follows by the follwing computation:
$(\nabla_{i}\otimes\nabla_{h_{0}}^{*})\Psi_{i}$ $=\nabla_{i}\circ\Psi_{i}-\Psi_{i}\circ\nabla_{h_{0}}$
$=\Psi_{i}\circ(\nabla_{h_{i}}-\nabla_{h_{0}})$
$=\Psi_{i}\circ(\Psi_{i}^{-2}\partial^{\nabla_{h_{0}}}(\Psi_{i}^{2}))$
$=2\partial^{\nabla_{h_{0}}}\Psi_{i}.$
3.3 Existence of $T$-adapted HYM metrics
Although this subsection overlaps with [Jin25-2, §4.2], we include the proof for the completeness.
Notation 3.16.
We use the notations in Notation 3.2. Let $Z=\overline{D\cup\Sigma}\subset X$ be the compact subset of Hausdorff codimension 2 and $\nabla_{\infty}$ be the $T$-admissible HYM connection of the complex hermitian vector bundle $(E,h_{0})|_{X\setminus Z}$ in Lemma 3.5. Let ${\Psi_{\infty}}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}:=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ and $\Phi_{\infty}:E_{\infty}\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ be holomorphic morphisms constructed in Theorem 3.13 and Theorem 3.14.
Theorem 3.17.
We use notations in Notation 3.16. Then the following holds.
  1. $(1)$
    $\Psi_{\infty}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}$ and $\Phi_{\infty}:E_{\infty}\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ satisfies $\Psi_{\infty}\circ\Phi_{\infty}={\rm{id}}$ and $\Phi_{\infty}\circ\Psi_{\infty}={\rm{id}}$. That is, $\Phi_{\infty}=\Psi_{\infty}^{-1}$. In particular $\Psi_{\infty}$ is a holomorphic isomorphism.
  2. $(2)$
    A smooth hermitian metric $h_{\infty}:=h_{0}{\Psi_{\infty}}^{2}$ on $E|_{X\setminus Z}$ is a $T$-adapted HYM metric on $E$ whose Chern connection is ${\Psi_{\infty}}^{-1}\circ\nabla_{\infty}\circ{\Psi_{\infty}}$.
Proof.
(1) By Theorem 3.13, we have $\lim_{i\to\infty}\Psi_{i}=\Psi_{\infty}$ almost everywhere on $X\setminus Z$, and by Theorem 3.14, we have $\lim_{i\to\infty}\Psi_{i}^{-1}=\Phi_{\infty}$ almost everywhere on $X\setminus Z$. Since $\Psi_{\infty}$ and $\Phi_{\infty}$ are both holomorphic, we obtain $\Psi_{\infty}\circ\Phi_{\infty}={\rm{id}}$ and $\Phi_{\infty}\circ\Psi_{\infty}={\rm{id}}$ on $X\setminus Z$ and thus $\Psi_{\infty}$ is a holomorphic isomorphism.
(2) By direct computation, we can see that ${\Psi_{\infty}}^{-1}\circ\nabla_{\infty}\circ{\Psi_{\infty}}$ is the Chern connection of $h_{\infty}$. By Theorem 3.13, it suffices to prove that the curvature tensor $F_{h_{\infty}}$ is $L^{2}$-integrable with respect to the norm defined by $h_{\infty}$ and $T$. We remark that $F_{h_{\infty}}={\Psi_{\infty}}^{-1}\circ F_{\nabla_{\infty}}\circ{\Psi_{\infty}}$ and $(\sqrt{-1}F_{h_{\infty}})^{*_{h_{\infty}}}=\sqrt{-1}F_{h_{\infty}}$. Then, the point-wise identity $|F_{h_{\infty}}|_{h_{\infty},T}^{2}T^{n}={\rm{Tr}}(\sqrt{-1}F_{h_{\infty}}\wedge\sqrt{-1}F_{h_{\infty}})\wedge T^{n}+|\Lambda_{T}F_{h_{\infty}}|_{h_{\infty}}^{2}T^{n}$ on $X\setminus Z$ (c.f. [Kob87] Chapter 4 section 4) shows $|F_{h_{\infty}}|_{h_{\infty},T}^{2}=|F_{\nabla_{\infty}}|_{h_{0},T}^{2}$ on $X\setminus Z$. Thus $h_{\infty}$ is a $T$-adapted HYM metric on $E|_{X\setminus Z}$. ∎
In conclusion, we obtain the following theorem.
Theorem 3.18.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $T$ be an adapted closed positive $(1,1)$-current in $\alpha$. Let $E$ be a holomorphic vector bundle on $X$. If $E$ is $\alpha^{n-1}$-slope stable, then $E$ admits a $T$-adapted HYM metric $h=h_{0}\Psi^{2}$ such that the following holds:
  1. $(1)$
    $\sup_{X\setminus D}|s_{D}\Psi|\leq C.$
  2. $(2)$
    $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi|^{2}T^{n}\leq C.$
Lemma 3.19.
We use notations in Notation 3.16. Let $h_{\infty}:=h_{0}\Psi_{\infty}^{2}$ be the $T$-adapted HYM metric on $E$ constructed in Theorem 3.17. Then the $T$-HYM constant $\lambda$ of $h_{\infty}$ satisfies the following:
$\lambda=\frac{1}{\alpha^{n}}\frac{\int_{X\setminus D}\sqrt{-1}{\rm{Tr}}F_{h_{\infty}}\wedge T^{n-1}}{{\rm{rk}}E}=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
Proof.
The first equality in the statement directly follows from the $T$-HYM equation of $h_{\infty}$. We show the second equation. Let $\nabla_{i}$ be the $\omega_{i}$-HYM connection on $(E,h_{0})$ in Lemma 3.5 and $\nabla_{\infty}:=\Psi_{\infty}\circ\nabla_{h_{\infty}}\circ\Psi_{\infty}^{-1}$ be the $T$-admissible HYM connection on $(E,h_{0})$ such that $\nabla_{i}\to\nabla_{\infty}$ in $C^{\infty}_{{\rm{loc}}}(X\setminus Z)$ as in Lemma 3.5. It is obvious that the $T$-HYM constant of $\nabla_{\infty}$ coincides with that of $h_{\infty}$. Since $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}\to\sqrt{-1}\Lambda_{T}F_{\nabla_{\infty}}$ at each point in $X\setminus D$ and since $\lambda_{i}$, the $\omega_{i}$-HYM constant of $\nabla_{i}$, satisfies $\lambda_{i}=\frac{1}{(\alpha+(1/i)\omega_{0})^{n}}\frac{c_{1}(E)\cdot(\alpha+(1/i)\omega_{0})^{n-1}}{{\rm{rk}}E}$, we obtain
$\lambda=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
3.4 Uniqueness and HYM $\Rightarrow$ stability
Notation 3.20.
  • Let $X$ be a compact Kähler manifold with a smooth Kähler metric $\omega_{0}$ on $X$ and $\alpha$ be a nef and big class on $X$. In this subsection, we use the following notations.
  • Assume that $D=E_{nK}(\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D,0}$ a smooth hermitian metric on $D$ with $|s_{D}|_{h_{D,0}}\leq 1$.
  • Let $T$ be a closed positive $(1,1)$-current in $\alpha$ satisfying $|s_{D}|^{k}\langle T^{n}\rangle\leq C\omega_{0}^{n}$ for some $k>0$ (refer to Lemma 2.17). Let $m\in\mathbb{Z}_{\geq 0}$ be a nonnegative integer so that $T\geq|s_{D}|^{2m}\omega_{0}$ holds.
  • Let $(E,h_{0})$ be a pair consisting of a holomorphic vector bundle $E$ over $X$ and a smooth hermitian metric $h_{0}$ on $E$.
We will use the following lemma that is shown by the standard computations.
Lemma 3.21.
We use the notations in Notation 3.20. Let $h=h_{0}\Psi$ be a $T$-adapted HYM metric on $E$ with $\det\Psi=1$. Assume that $\sup_{X\setminus D}|s_{D}^{m}\Psi|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}^{m}}\leq C$ and
$\int_{X\setminus D}|s_{D}^{m}\partial^{\nabla_{h_{0}}}\Psi|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0},T}^{2}T^{n}\leq C$. Then, for the inverse morphism $\Psi^{-1}$ of $\Psi$, the following holds:
  • $\sup_{X\setminus D}|s_{D}^{m}\Psi^{-1}|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}\leq C$.
  • $\int_{X\setminus D}|s_{D}^{3m}\partial^{\nabla_{h_{0}}}\Psi^{-1}|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0^{3m}}}^{2}T^{n}\leq C$.
We firstly establish the following lemma.
Lemma 3.22.
We use notations in Notation 3.20. Fix $l>0$ sufficently large. Let $h_{D}:=h_{D,0}\Psi_{D}^{2}$ be a $T$-adapted HYM metric on $\mathcal{O}(lD)$ constructed in Theorem 3.17. Let us consider the following composition of holomorphic morphisms:
$(\mathcal{O}_{X},h_{{\rm triv}})\xrightarrow{s_{D}^{l}}(\mathcal{O}(lD),h_{D,0})\xrightarrow{\Psi_{D}}(\mathcal{O}(lD)|_{X\setminus D},\nabla_{D})$
where $h_{{\rm triv}}$ is the trivial metric on $\mathcal{O}_{X}$ and $\nabla_{D}=\Psi_{D}^{-1}\circ\nabla_{h_{D}}\circ\Psi_{D}$ is the $T$-admissible HYM connection on $(\mathcal{O}(lD),h_{D,0})|_{X\setminus D}$. Then the following holds:
[…]
Proof.
We recall that, by the construction of $\Psi_{D}$ (see Theorem 3.13 and Proposition 3.15), it satisies
$\sup_{X\setminus D}|s_{D}\Psi_{D}|\leq C\hbox{ and }\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{D,0}}}\Psi_{D}|^{2}\leq C.$ (3.35)
Since $s_{D}^{l}:(\mathcal{O}_{X},\bar{\partial}^{\nabla_{h_{{\rm triv}}}})\to(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})$ and $\Psi_{D}:(\mathcal{O}(lD),\bar{\partial}^{\nabla_{h_{D,0}}})\to(\mathcal{O}(lD)|_{X\setminus D},\bar{\partial}^{\nabla_{D}})$ is holomorphic (refer to Theorem 3.13) and since $\sqrt{-1}\Lambda_{T}F_{h_{D}}=0$ by Lemma 3.19, we obtain
$\Delta_{T}|\Psi_{D}s_{D}^{l}|_{h_{D,0}\otimes h_{{\rm triv}}}^{2}=\left|(\nabla_{D}\otimes\nabla_{h_{{\rm triv}}}^{*})(\Psi_{D}s_{D}^{l})\right|^{2}_{h_{D,0}\otimes h_{{\rm triv}}^{*},T}.$ (3.36)
We show that the integral of LHS with respect to $T^{n}$ equals to 0. Let $\eta_{k}:X\to[0,1]$ be a smooth cut-off functions as in Lemma 3.2, that is, each $\eta_{k}$ satisfies
${\rm{Supp}}(\eta_{k})\Subset X\setminus D,\hskip 5.69054ptX\setminus D_{\varepsilon_{k-1}}\subset\{\eta_{k}=1\},\hskip 5.69054pt\lim_{k\to\infty}\int_{X}|d\eta_{k}|^{2}_{\omega_{0}}\omega_{0}^{n}=0.$
Then we comute as follows:
$\left|\nabla|\Psi_{D}s_{D}^{l}|^{2}\right|_{T}$ $\leq 2\left|\langle(\nabla_{D}\otimes\nabla_{h_{D,0}}^{*}\Psi_{D})s_{D}^{l}+l\Psi_{D}s_{D}^{l-1}\nabla_{h_{D,0}}s_{D},\Psi_{D}s_{D}^{l}\rangle\right|$
$\leq|s_{D}|^{m+1}|\partial^{\nabla_{h_{D,0}}}\Psi_{D}|_{T}|s_{D}\Psi_{D}||s_{D}|^{2l-m-2}+l|s_{D}\Psi_{D}|^{2}|s_{D}|^{2m}|\nabla_{h_{D,0}}s_{D}|_{T}|s_{D}|^{2l-2m-3}$
$\leq C\left(|s_{D}^{m+1}\partial^{\nabla_{h_{D,0}}}\Psi_{D}|_{T}+1\right)|s_{D}|^{2l-2m-3}$
In the last line, we used (3.35) and the assumption that $T\geq|s_{D}|^{2m}\omega_{0}$. Then we can estimate as follows:
$\left|\int_{X\setminus D}\eta_{k}\Delta_{T}|\Psi_{D}s_{D}^{l}|^{2}T^{n}\right|$ $=\left|\int_{X\setminus D}\langle\nabla\eta_{k},\nabla|\Psi_{D}s_{D}^{l}|^{2}\rangle_{T}T^{n}\right|$
$\leq\int_{X\setminus D}|\nabla\eta_{k}|_{T}|\nabla|\Psi_{D}s_{D}^{l}|^{2}|_{T}T^{n}$
$\leq C\left(\int_{X}|\nabla\eta_{k}|_{\omega_{0}}^{2}|s_{D}|^{4l-2m-6}T^{n}\right)^{1/2}\left(\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{D,0}}}\Psi_{D}|_{T}^{2}T^{n}\right)^{1/2}$
$\hskip 11.38109pt+C\int_{X}|\nabla\eta_{k}|_{\omega_{0}}|s_{D}|^{2l-4m-3}T^{n}$
$\leq C\left(\int_{X}|\nabla\eta_{k}|_{\omega_{0}}^{2}\omega_{0}^{n}\right)^{1/2}\left(\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{D,0}}}\Psi_{D}|_{T}^{2}T^{n}\right)^{1/2}$
$\hskip 11.38109pt+C\int_{X}|\nabla\eta_{k}|_{\omega_{0}}\omega_{0}^{n}$
In the last line, we use $|s_{D}|^{k}T^{n}\leq C\omega_{0}^{n}$ (see Lemma 2.17). Hence we obtain
$\int_{X\setminus D}\Delta_{T}|\Psi_{D}s_{D}^{l}|^{2}T^{n}=\lim_{k\to\infty}\int_{X\setminus D}\eta_{k}\Delta_{T}|\Psi_{D}s_{D}^{l}|^{2}T^{n}=0.$ (3.37)
Then, by (3.36) and (3.37), we obtain that
[…]
Theorem 3.23.
If $(E,h_{0})$ admits two $T$-adapted HYM metric $h_{1}$ and $h_{2}$, then the Chern connections $\nabla_{h_{1}}$ of $h_{1}$ coincides with the Chern connection $\nabla_{h_{2}}$ of $h_{2}$, that is, we have $\nabla_{h_{1}}=\nabla_{h_{2}}$.
Proof.
Let $h_{1}=h_{0}\Psi_{1}^{2}$ and $h_{2}=h_{0}\Psi_{2}^{2}$ be two $T$-adapted HYM metrics with $\det\Psi_{i}=1$ for $i=1,2$. Suppose that the $T$-HYM constants $\lambda_{i}$ of $h_{i}$ satisfies $\lambda_{1}\geq\lambda_{2}$. Let us consider a connection $\nabla_{i}:=\Psi_{i}\circ\nabla_{h_{i}}\circ\Psi_{i}^{-1}$ on $E$. Then $\nabla_{i}$ is a $T$-admissible HYM connection of the complex hermitian vector bundle $(E,h_{0})$ and
$\Psi_{i}:(E,\bar{\partial}^{E})|_{X\setminus D}\to(E|_{X\setminus D},\bar{\partial}^{\nabla_{i}})$
is a holomorphic isomorphism satisfying, for $m$ in Notation 3.20,
$\sup_{X\setminus D}|s_{D}^{l}\Psi_{i}|_{h_{0}}\leq C\hbox{ for some $l\in\mathbb{Z}_{\geq m}$}$ (3.38)
and
$\int_{X\setminus D}\left|(\nabla_{i}\otimes\nabla_{h_{0}}^{*})(s_{D}^{l}\Psi_{i})\right|^{2}_{h_{0}^{*}\otimes h_{0}\otimes h_{D,0}^{l},T}T^{n}\leq C.$ (3.39)
By Theorem 3.17, the holomorphic line bundle $\mathcal{O}((4l+1)D)$ admits a $T$-adapted HYM metric $h_{D}=h_{D,0}\Psi_{D}^{2}$ with $\det\Psi_{D}=1$. Let us define $\nabla_{D}:=\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1}$ a $T$-adapted HYM connection of $(\mathcal{O}((4l+1)D),h_{D,0})$. Then, by Theorem 3.13 and Proposition 3.15, we have that $\Psi_{D}$ satisfies
$\sup_{X\setminus D}|s_{D}\Psi_{D}|_{h_{D,0}^{4l+1,*}\otimes h_{D,0}^{4l+1}\otimes h_{D,0}}$ (3.40)
and
$\int_{X\setminus D}\left|s_{D}^{m+1}\partial^{\nabla_{h_{D,0}}}\Psi_{D}\right|^{2}_{h_{D,0}^{*}\otimes h_{D,0}\otimes h_{D,0}^{m+1},T}T^{n}\leq C.$ (3.41)
Let us consider the following diagram. All vector bundles are restricted to $X\setminus D$:
$(E,\bar{\partial}^{E},\nabla_{h_{0}},h_{0})$$(E,\bar{\partial}^{\nabla_{1}},\nabla_{1},h_{0})$$(E,\bar{\partial}^{\nabla_{2}},\nabla_{2},h_{0})$$(E\otimes\mathcal{O}((4l+1)D),\bar{\partial}^{\nabla_{2}\otimes\nabla_{h_{D,0}}},\nabla_{2}\otimes\nabla_{h_{D,0}},h_{0}\otimes h_{D,0})$$(E\otimes\mathcal{O}((4l+1)D),\bar{\partial}^{\nabla_{2}\otimes\nabla_{D}},\nabla_{2}\otimes\nabla_{D},h_{0}\otimes h_{D,0})$$\Psi_{1}^{-1}$$\Psi_{2}$${\rm{id}}_{E}\otimes s_{D}^{4l+1}$${\rm{id}}_{E}\otimes\Psi_{D}$$s$
Here $s:=({\rm{id}}_{E}\otimes\Psi_{D})\circ({\rm{id}}_{E}\otimes s_{D}^{4l+1})\circ\Psi_{2}\circ\Psi_{1}^{-1}$ is a holomorphic morphism from $(E,\bar{\partial}^{\nabla_{1}})$ to $(E\otimes\mathcal{O}((4l+1)D),\bar{\partial}^{\nabla_{2}\otimes\nabla_{D}})$. We remark that the $T$-HYM constant of $\nabla_{D}$ equals to 0 by Lemma 3.19. Then the usual calculation shows the following point-wise equation on $X\setminus D$:
$\Delta_{T}|s|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}=\left|(\nabla_{2}\otimes\nabla_{1}^{*}\otimes\nabla_{D})s\right|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0},T}-(\lambda_{2}-\lambda_{1})|s|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}.$ (3.42)
By (3.38), (3.39), (3.40) and (3.41), we can see that each term in RHS in (3.42) is integrable over $X\setminus D$ with respect to $T^{n}$. Let $\eta_{k}$ be smooth cut-off functions on $X$ as in Lemma 3.2, in particular they satisfies ${\rm{Supp}}(\eta_{k})\Subset X\setminus D$, $\eta_{k}\nearrow 1$ and $\int_{X}|d\eta_{k}|^{2}_{\omega_{0}}\omega_{0}^{n}\to 0$. Then, we obtain the following equation by the same way with (3.37):
$\int_{X\setminus D}\Delta_{T}|s|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}^{2}T^{n}=0.$ (3.43)
Here we recall that the assumption $\lambda_{1}\geq\lambda_{2}$. Then, by integrating both sides of (3.42), we obtain
$\lambda_{1}=\lambda_{2}\hbox{ and }\left|(\nabla_{2}\otimes\nabla_{1}^{*}\otimes\nabla_{D})s\right|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0},T}=0.$
In particular, we have
$0$ $=\left(\nabla_{2}\otimes\nabla_{1}^{*}\otimes\nabla_{D}\right)s$
$=\left(\nabla_{2}\otimes\nabla_{1}^{*}\otimes\nabla_{D}\right)(s_{D}^{4l+1}\circ\Psi_{D}\circ\Psi_{2}\circ\Psi_{1}^{-1})$
$=(\nabla_{D}\otimes\nabla_{0}^{*})(s_{D}^{4l+1}\Psi_{D})\circ\Psi_{2}\circ\Psi_{1}^{-1}+s_{D}^{4l+1}\Psi_{D}\circ(\nabla_{2}\otimes\nabla_{1}^{*})(\Psi_{2}\circ\Psi_{1}^{-1}).$ (3.44)
By Lemma 3.22, we have $(\nabla_{D}\otimes\nabla_{0}^{*})(s_{D}^{4l+1}\Psi_{D})=0$. Then, by (3.4), we have
$(\nabla_{2}\otimes\nabla_{1}^{*})(\Psi_{2}\circ\Psi_{1}^{-1})=0.$ (3.45)
We remark that
$|\Psi_{2}\circ\Psi_{1}^{-1}|^{2}_{h_{0}\otimes h_{0}^{*}}=|{\rm{id}}_{E}|^{2}_{h_{2}\otimes h_{1}^{*}}$ (3.46)
and
$\Delta_{T}|{\rm{id}}_{E}|^{2}_{h_{2}\otimes h_{1}^{*}}=|(\nabla_{h_{2}}\otimes\nabla_{h_{1}}^{*})({\rm{id}}_{E})|^{2}_{h_{2}\otimes h_{1}^{*},T}-(\lambda_{2}-\lambda_{1})|{\rm{id}}_{E}|^{2}_{h_{2}\otimes h_{1}^{*}},$ (3.47)
furthermore
$\Delta_{T}|\Psi_{2}\circ\Psi_{1}^{-1}|_{h_{0}\otimes h_{0}^{*}}=|(\nabla_{2}\otimes\nabla_{1}^{*})(\Psi_{2}\circ\Psi_{1}^{-1})|_{h_{0}\otimes h_{0}^{*},T}-(\lambda_{2}-\lambda_{1})|\Psi_{2}\circ\Psi_{1}^{-1}|_{h_{0}\otimes h_{0}^{*}}$ (3.48)
at each point of $X\setminus D$. Then by (3.46), (3.47) and (3.48) together with (3.45), we obtain
$|(\nabla_{h_{2}}\otimes\nabla_{h_{1}}^{*}){\rm{id}}_{E}|^{2}_{h_{2}\otimes h_{1}^{*},T}=\left|(\nabla_{2}\otimes\nabla_{1}^{*})(\Psi_{2}\circ\Psi_{1}^{-1})\right|^{2}_{h_{0}\otimes h_{0}^{*},T}=0.$ (3.49)
The equation (3.49) shows $\nabla_{h_{1}}=\nabla_{h_{2}}$. ∎
Corollary 3.24.
We use the notations in Notation 3.20. Assume that a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric $h$. Then $h$ computes the $\alpha^{n-1}$-slope of $E$:
$\mu_{\alpha}(E)=\frac{1}{{\rm{rk}}E}\int_{X\setminus D}c_{1}(E,h)\wedge\frac{T^{n-1}}{(n-1)!}$
where $c_{1}(E,h)=\sqrt{-1}Tr(F_{h})$. In particular, the $T$-HYM constant $\lambda$ of $h$ is described as follows:
$\lambda=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
Proof.
Let $h$ be the $T$-adapted HYM metric on $E$. By Lemma 3.5 and Theorem 3.17, the complex hermitian line bundle $(\det E,\det h_{0})$ admits the $\omega_{i}$-HYM connection $\nabla_{i}$ and the $T$-adapted HYM connection $\nabla_{\det E}$ such that $\nabla_{i}\to\nabla_{\det E}$ locally smoothly on $X\setminus D$. We furthermore have the holomorphic isomorphism $\Psi_{\det E}:(\det E,\bar{\partial}^{E})|_{X\setminus D}\to(\det E|_{X\setminus D},\bar{\partial}^{\nabla_{\det E}})$ appeared in Theorem 3.17. Since $\nabla_{i}$ is $\omega_{i}$-HYM, it satisfies the $\omega_{i}$-HYM equation $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}=\lambda_{i}{\rm{id}}$ on $X\setminus D$ with
$\lambda_{i}=\frac{c_{1}(\det E)\cdot\{\omega_{i}\}^{n-1}}{\{\omega_{i}\}^{n}}.$
Since $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}\to\sqrt{-1}\Lambda_{T}F_{\nabla_{\det E}}$ at each point in $X\setminus D$, we obtain
$\sqrt{-1}\Lambda_{T}F_{\nabla_{\det E}}=\lambda\hbox{ with $\lambda=\frac{c_{1}(\det E)\cdot\alpha^{n-1}}{\alpha^{n}}.$}$ (3.50)
By Theorem 3.17, the pull-back metric $\Psi_{\det E}^{*}(\det h_{0})$ is a $T$-adapted HYM metric on $\det E$ with the chern connection $\Psi_{\det E}^{-1}\circ\nabla_{\det E}\circ\Psi_{\det E}$. By the assumption, we also have a $T$-adapted HYM metric $\det h$ on $\det E$ induced by the $T$-adapted HYM metric $h$ on $E$. By Theorem 3.23, we have $\Psi_{\det E}^{-1}\circ\nabla_{\det E}\circ\Psi_{\det E}=\nabla_{\det h}$. Then, by (3.50), we obtain
${\rm{Tr}}(\sqrt{-1}\Lambda_{T}F_{h})=\lambda.$
Let $\widetilde{\lambda}$ be the $T$-HYM constant of $h$, that is, it satisfies
$\sqrt{-1}\Lambda_{T}F_{h}=\widetilde{\lambda}{\rm{id}}_{E}.$
Then, by taking trace and integrating over $X\setminus D$ with respect to $T^{n}$, we obtain
${\rm{rk}}E\alpha^{n}\widetilde{\lambda}=\int_{X\setminus D}{\rm{Tr}}(\sqrt{-1}\Lambda_{T}F_{h})T^{n}=\alpha^{n}\lambda.$
Hence, together with (3.50), we obtain the result.
We need the following lemma to prove the uniqueness of $T$-adapted HYM metrics.
Lemma 3.25.
We use the notations in Notation 3.20. Assume that holomorphic vector bundles $E_{1}$ and $E_{2}$ are $\alpha^{n-1}$-slope stable with ${\rm{rk}}E_{1}={\rm{rk}}E_{2}$ and $\mu_{\alpha}(E_{1})=\mu_{\alpha}(E_{2})$. Let $g:E_{1}|_{X\setminus D}\to E_{2}|_{X\setminus D}$ be a nonzero holomorphic morphism satisfying
$\sup_{X\setminus D}|s_{D}^{l}g|_{h_{2}\otimes h_{1}^{*}\otimes h_{D,0}}<\infty$
for some $l\in\mathbb{Z}_{\geq 0}$ where $h_{i}$ is a smooth hermitian metric on $E_{i}$ for $i=1,2$. Then $g$ is isomorphic. In particular, if $E_{1}=E_{2}=E$, then there is a constant $a\in\mathbb{C}\setminus\{0\}$ such that $g=a{\rm{id}}_{E}$ on $X\setminus D$.
Proof.
Let us consider the image sheaf $\mathcal{G}$ of $s_{D}^{l}g$, which is a coherent subsheaf of $E_{2}\otimes\mathcal{O}(lD)|_{X\setminus D}$. Since $s_{D}^{l}g$ is bounded and $E_{1}$ is a locally free sheaf defined all over $X$, the coherent sheaf $\mathcal{G}$ on $X\setminus D$ is generated by bounded sections of $E_{2}\otimes\mathcal{O}(lD)|_{X\setminus D}$ around $D$. Then, since $E_{2}\otimes\mathcal{O}(lD)$ is defined all over $X$, bounded generators of $\mathcal{G}$ extends across $D$ as a sections of $E_{2}\otimes\mathcal{O}(lD)|_{X\setminus D}$. Hence, the image sheaf $\mathcal{G}$ extends to a coherent subsheaf of $E_{2}\otimes\mathcal{O}(lD)$, also denoted by $\mathcal{G}$. We remark that $s_{D}^{l}g\in H^{0}(X\setminus D,E_{2}\otimes E_{1}^{*}\otimes\mathcal{O}(lD))$ also extends to a holomorphic morphism defined all over $X$, since $s_{D}^{l}g$ is bounded and $E_{1}$ and $E_{2}\otimes\mathcal{O}(lD)$ are both defined over $X$. Then we obtain the following short exact sequence of coherent sheaves on $X$:
$0\rightarrow\mathcal{F}\rightarrow E_{1}\xrightarrow{s_{D}^{l}g}\mathcal{G}\rightarrow 0.$
Then, since $E_{1}$ and $E_{2}$ are $\alpha^{n-1}$-slope stable with same $\alpha^{n-1}$-slope, we can see that $\mathcal{F}=0$. In fact, the $\alpha^{n-1}$-slope stability of $E_{2}$ ensures that $E_{2}\otimes\mathcal{O}(lD)$ is also $\alpha^{n-1}$-slope stable. We recall that $\mathcal{G}\subset E_{2}\otimes\mathcal{O}(lD)$. Then if $\mathcal{F}\neq 0$, we have $\mu_{\alpha}(E_{1})<\mu_{\alpha}(\mathcal{G})<\mu_{\alpha}(E_{2}\otimes\mathcal{O}(lD))=\mu_{\alpha}(E_{1})$, that is a contradiction. Therefore $s_{D}^{l}g:E_{1}\to E_{2}\otimes\mathcal{O}(lD)$ is injective. Hence its determinant gives the short exact sequence
$0\to\det E_{1}\xrightarrow{\det(s_{D}^{l}g)}\det(E_{2}\otimes\mathcal{O}(lD))\to\tau_{V}\to 0.$
Since $s_{D}^{l}g$ is injective, the quotient sheaf $\tau_{V}$ is a torsion sheaf supported on some analytic subvariety $V\subset X$. We recall the assumptions that ${\rm{rk}}E_{1}={\rm{rk}}E_{2}$ and $\mu_{\alpha}(E_{1})=\mu_{\alpha}(E_{2})$. Then we have
$0=c_{1}(\det(E_{2}\otimes\mathcal{O}(lD)))\alpha^{n-1}-c_{1}(\det E_{1})\alpha^{n-1}=c_{1}(\tau_{V})\alpha^{n-1}.$
Hence we obtain a decomposition $V=V_{1}\cup V_{2}$ such that $V_{1}\subset D$ and ${\rm{codim}}V_{2}\geq 2$. In particular we have that $\det(s_{D}^{l}g)$ is isomorphic over $X\setminus D$ and thus it has no nontrivial zero locus in $X\setminus D$. As a consequence $g:E_{1}|_{X\setminus D}\to E_{2}|_{X\setminus D}$ is isomorphic. Assume that $E=E_{1}=E_{2}$ and let $a\in\mathbb{C}\setminus\{0\}$ be an eigenvalue of $g$ at $x\in X\setminus D$. If $g-a{\rm{id}}_{E}\neq 0$, then it is an isomorphism by the above proof. But it is a contradiction. Hence $g=a{\rm{id}}_{E}$. ∎
Then we can prove the uniqueness of a $T$-adapted HYM metric on an $\alpha^{n-1}$-slope stable vector bundle.
Lemma 3.26.
We use the notations in Notation 3.20. If $E$ is $\alpha^{n-1}$-slope stable, then a $T$-adapted HYM metric on $E$ is unique up to scaling, if it exists.
Proof.
Let $h_{1}$ and $h_{2}$ be two $T$-adapted HYM metrics on a holomorphic vector bundle $E$. We denote by $h_{2}=h_{1}g$ for some positive definite $h_{1}$-hermitian endomorphism $g$. Let $l\in\mathbb{Z}_{>0}$ be a sufficiently large integer. Let $h_{D}$ be a $T$-adapted HYM metric on $\mathcal{O}(lD)$ whose existence is proved in Theorem 3.17. Let us define a smooth section $\overline{s_{D}^{l}}:=h_{D}(\cdot,s_{D}^{l})$ of $\mathcal{O}(lD)^{*}=\mathcal{H}om(\mathcal{O},\mathcal{O}(lD)^{*})$ over $X\setminus D$. If we denote by $h_{D}=h_{D,0}\Psi_{D}^{2}$, then $\overline{s_{D}^{l}}=h_{D,0}(\Psi_{D}\circ\cdot,\Psi_{D}s_{D}^{l})$ and we have, by applying Lemma 3.22 in the first equality below,
$0$ $=(\nabla_{D}\otimes\nabla_{triv}^{*})(\Psi_{D}s_{D}^{l})$
$=(\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1})\circ\Psi_{D}s_{D}^{l}-\Psi_{D}s_{D}^{l}\circ\nabla_{triv}$
$=\Psi_{D}\circ(\nabla_{h_{D}}\otimes\nabla_{triv}^{*})s_{D}^{l}$
$=\Psi_{D}\circ(\nabla_{h_{D}}s_{D}^{l})$
Here $\nabla_{D}=\Psi_{D}\circ\nabla_{h_{D}}\circ\Psi_{D}^{-1}$ the $T$-adapted HYM connection on $(\mathcal{O}(lD),h_{D,0})$ and $\nabla_{triv}$ is the trivial connection on the trivial bundle $\mathcal{O}$. Since $\Psi_{D}$ does not have zero locus on $X\setminus D$, we obtain
$\nabla_{h_{D}}s_{D}^{l}=0.$ (3.51)
Then we also obtain that for any smooth section $t$ of $\mathcal{O}(lD)$,
$(\nabla_{h_{D}}\overline{s_{D}^{l}})(t)$ $=d\left(h_{D}(t,s_{D}^{l})\right)-h_{D}(\nabla_{h_{D}}t,s_{D}^{l})$
$=-h_{D}(t,\nabla_{h_{D}}s_{D}^{l})=0$ (3.52)
by (3.51). Then, we can see the following:
Claim 3.27.
$\bar{\partial}^{E\otimes E^{*}\otimes\mathcal{O}(lD)^{*}}\left(\overline{s_{D}^{l}}g\right)=0$. In particular, we have
$\bar{\partial}^{E\otimes E^{*}}g=0.$
proof of the claim.
In fact, by the uniqueness of a $T$-adapted HYM connection of $E$ (Theorem 3.23), we have
$0=\nabla_{h_{1}}-\nabla_{h_{2}}=g^{-1}\partial^{\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}}g.$
In particular we have
$\partial^{\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}}g=0$ (3.53)
on $X\setminus D$ since $g$ is positive definite. If we denote the $T$-adapted HYM metric $h_{D}$ on $\mathcal{O}(lD)$ by $h_{D}=h_{D,0}\Psi_{D}^{2}$, then we can assume $|s_{D}\Psi_{D}|_{h_{D,0}\otimes h_{D,0}^{*}}$ is bounded by Theorem 3.13 (2). If we further remark that $h_{1}$ and $h_{2}$ are both $T$-adapted, for sufficiently large $l\in\mathbb{Z}_{>0}$ we have
$\sup_{X\setminus D}|\overline{s_{D}^{l}}g|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}<\infty,$ (3.54)
$\sup_{X\setminus D}|\overline{s_{D}^{l}}g|_{h_{1}\otimes h_{1}^{*}\otimes h_{D}}<\infty,$ (3.55)
$\int_{X\setminus D}\left|(\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}\otimes\nabla_{h_{D}})\overline{s_{D}^{l}}g\right|_{h_{1}\otimes h_{1}^{*}\otimes h_{D},T}T^{n}<\infty.$ (3.56)
If we use $\partial^{\nabla_{h_{D}}}\overline{s_{D}^{l}}=0$ (3.4) together with (3.53), we have
$\partial^{\nabla_{h_{1}}\otimes\nabla_{h_{1}}^{*}\otimes\nabla_{h_{D}}}(\overline{s_{D}^{l}}g)=0.$ (3.57)
Thus we obtain
$\Delta_{T}|\overline{s_{D}^{l}}g|^{2}_{h_{1}\otimes h_{1}^{*}\otimes h_{D}}=-2|\bar{\partial}^{E\otimes E^{*}\otimes\mathcal{O}(lD)^{*}}\overline{s_{D}^{l}}g|_{h_{1}\otimes h_{1}^{*}\otimes h_{D},T}^{2}.$ (3.58)
by a standard computation. By (3.55) and (3.56) we can see that the both sides of (3.58) are integrable over $X\setminus D$ with respect to $T^{n}$ and by the same computation with (3.37) we obtain
$0=\int_{X\setminus D}\Delta_{T}|\overline{s_{D}^{l}}g|^{2}_{h_{1}\otimes h_{1}^{*}\otimes h_{D}}T^{n}=-2\int_{X\setminus D}|\bar{\partial}^{E\otimes E^{*}\otimes\mathcal{O}(lD)^{*}}\overline{s_{D}^{l}}g|_{h_{1}\otimes h_{1}^{*}\otimes h_{D},T}^{2}T^{n}.$
Hence we have
$\bar{\partial}^{E\otimes E^{*}\otimes\mathcal{O}(lD)^{*}}\left(\overline{s_{D}^{l}}g\right)=0.$ (3.59)
Then, by (3.59) and and (3.4), we obtain that
$\bar{\partial}^{E\otimes E^{*}}g=0.$
The proof of Claim 3.27 ends. ∎
Since $h_{1}$ and $h_{2}$ are $T$-adapted and $g$ is defined by $h_{2}=h_{1}g$, there is $l\in\mathbb{Z}_{\geq 0}$ such that $\sup_{X\setminus D}|s_{D}^{l}g|_{h_{0}\otimes h_{0}^{*}\otimes h_{D,0}}<\infty$. Then, by Claim 3.27 and Lemma 3.25, we can conclude that $g=a{\rm{id}}_{E}$ over $X\setminus D$ for some $a\in\mathbb{R}\setminus\{0\}$ since now we assume that $E$ is $\alpha^{n-1}$-slope stable. As a consequence, we obtain $h_{2}=ah_{1}$ and thus we obtain the uniqueness of $T$-adapted HYM metrics on $E$. ∎
Theorem 3.28.
We use the notations in Notation 3.20. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then $E$ is $\alpha^{n-1}$-slope polystable.
Proof.
We denote by $h_{E}$ the $T$-adapted HYM metric on $E$. Let $\iota:\mathcal{F}\hookrightarrow E$ be an $\alpha^{n-1}$-slope stable torsion free subsheaf of rank $q$ with torsion free quotient $\mathcal{G}:=E/\mathcal{F}$ such that
$\mu_{\alpha}(E)\leq\mu_{\alpha}(\mathcal{F}).$ (3.60)
Let $\pi:Y\to X$ be a sequence of blow-ups so that $F:=\pi^{*}\mathcal{F}/{\rm{Tor}}$ is locally free. Then we have $\mu_{\pi^{*}\alpha}(F)=\mu_{\alpha}(\mathcal{F})\geq\mu_{\alpha}(E)=\mu_{\pi^{*}\alpha}(\pi^{*}E)$. We denote by $i:F\to\pi^{*}E$ the induced morphism from the inclusion $\iota:\mathcal{F}\hookrightarrow E$. We remark that $\pi^{*}T$ is an adapted current on $Y$ (Lemma 2.19), and $F$ is $\pi^{*}\alpha^{n-1}$-slope stable (Lemma 2.12), $F$ admits a $T$-adapted HYM metric $h_{F}$. Let $D^{\prime}:=E_{nK}(\pi^{*}\alpha)$ be a snc divisor in $Y$ given by the non-Kähler locus of $\pi^{*}\alpha$. We denote by $h_{D^{\prime}}$ the $\pi^{*}T$-adapted HYM metric on $\mathcal{O}(lD)$ for sufficiently large $l\in\mathbb{Z}_{>0}$. Then we consider the following composition of morphisms of $\pi^{*}T$-adapted HYM vector bundles:
$(F,h_{F})\xrightarrow{i}(\pi^{*}E,\pi^{*}h_{E})\xrightarrow{s_{D^{\prime}}^{l}}(\pi^{*}E\otimes\mathcal{O}(lD^{\prime}),\pi^{*}h_{E}\otimes h_{D^{\prime}}).$
Then, since $s_{D^{\prime}}^{l}\circ i$ is holomorphic, we have
$\Delta_{T}|s_{D^{\prime}}^{l}i|^{2}_{\pi^{*}h_{E}\otimes h_{F}\otimes h_{D^{\prime}}}$
$=|(\nabla_{\pi^{*}h_{E}}\otimes\nabla_{h_{F}}\otimes\nabla_{h_{D^{\prime}}})s_{D^{\prime}}^{l}i|^{2}_{\pi^{*}h_{E}\otimes h_{F}\otimes h_{D^{\prime}},\pi^{*}T}-(\lambda_{E}-\lambda_{F})|s_{D^{\prime}}^{l}i|^{2}_{\pi^{*}h_{E}\otimes h_{F}\otimes h_{D^{\prime}}}$ (3.61)
where $\lambda_{E}$ is the $T$-HYM constant of $E$ and $\lambda_{F}$ is the $\pi^{*}T$-HYM constant of $F$. By Corollary 3.24 and (3.60), we have $\lambda_{E}\leq\lambda_{F}$. Since $\pi^{*}h_{E}$, $h_{F}$ and $h_{D^{\prime}}$ are all $\pi^{*}T$-adapted, each term in (3.4) is integrable over $Y\setminus D^{\prime}$ with respect to $T^{n}$ (refer to (3.42)). Furthermore, by the cut-off argument as (3.37), we obtain that
$\int_{Y\setminus D^{\prime}}\Delta_{T}|s_{D^{\prime}}^{l}i|^{2}_{\pi^{*}h_{E}\otimes h_{F}\otimes h_{D^{\prime}}}(\pi^{*}T)^{n}=0$
Then we obtain that $\lambda_{E}=\lambda_{F}$, in particular
$\mu_{\alpha}(E)=\mu_{\alpha}(\mathcal{F}),$ (3.62)
and
$0=(\nabla_{\pi^{*}h_{E}}\otimes\nabla_{h_{F}}\otimes\nabla_{h_{D^{\prime}}})s_{D^{\prime}}^{l}i=(\nabla_{h_{D^{\prime}}}s_{D^{\prime}}^{l})i+s_{D^{\prime}}^{l}(\nabla_{\pi^{*}h_{E}}\otimes\nabla_{h_{F}}^{*})i.$ (3.63)
By (3.51), we have $\nabla_{h_{D^{\prime}}}s_{D^{\prime}}^{l}=0$. Hence (3.63) implies
$(\nabla_{\pi^{*}h_{E}}\otimes\nabla_{h_{F}}^{*})i=0.$ (3.64)
Let us consider a short exact sequence
$0\to\mathcal{F}\xrightarrow{\iota}E\rightarrow\mathcal{G}\to 0$ (3.65)
over $X$ and a short exact sequence
$0\to F|_{Y\setminus D^{\prime}}\xrightarrow{i}\pi^{*}E|_{Y\setminus D^{\prime}}\rightarrow G\to 0$ (3.66)
over $Y\setminus D^{\prime}$. By (3.64), the sequence (3.66) holomorphically splits: $\pi^{*}E|_{Y\setminus D^{\prime}}\simeq F|_{Y\setminus D^{\prime}}\oplus G$. Since $F|_{Y\setminus D^{\prime}}=\pi^{*}(\mathcal{F}|_{X\setminus(D\cup{\rm{Sing}}\mathcal{F})})$ and $G=\pi^{*}(\mathcal{G}|_{X\setminus(D\cup{\rm{Sing}}\mathcal{F})})$, and the codimension of $\rm{Sing}$$\mathcal{F}$, the non locally free locus of $\mathcal{F}$, is at least 3, we obtain that the sequence (3.65) holomorphically splits: $E|_{X\setminus D}\simeq(\mathcal{F}\oplus\mathcal{G})|_{X\setminus D}$. By (3.62), we have $\mu_{\alpha}(E)=\mu_{\alpha}(\mathcal{G})$. We remark that $\mathcal{G}$ also admits a $T$-adapted HYM metric, since $E$ and $\mathcal{F}$ both admit. Then, by repeating this argument by replacing $E$ to $\mathcal{G}$, we obtain the $\alpha^{n-1}$-slope polystability of $E$. ∎
Then we can prove the uniqueness in the general setting.
Theorem 3.29.
We use the notations in Notation 3.20. If a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric, then it is unique up to scaling.
Proof.
Let $h_{1}$ and $h_{2}$ be two $T$-adapted HYM metrics on $E$. Then, by Theorem 3.28, we can see that $E$ is $\alpha^{n-1}$-slope polystable. That is, there exists $\alpha^{n-1}$-slope stable torsion free sheaves $\mathcal{E}_{1},\ldots,\mathcal{E}_{k}$ on $X$ such that $E$ holomorphically splits to the direct sum of $\mathcal{E}_{i}$ on $X\setminus D$, that is, there is a holomorphic isomorphism $E|_{X\setminus D}\simeq(\mathcal{E}_{1}\oplus\cdots\mathcal{E}_{k})|_{X\setminus D}$. By Theorem 3.26, the $T$-adapted HYM metric $g_{i}$ on $\mathcal{E}_{i}$ is unique. As in the proof of Theorem 3.28, we can assume that $\iota:\mathcal{E}_{1}\hookrightarrow E$ is an $\alpha^{n-1}$-stable subsheaf and we have by (3.64),
$(\nabla_{h_{i}}\otimes\nabla_{g_{1}}^{*})\iota=0$
on the locally free locus of $\mathcal{E}_{1}$ for $i=1,2$. Thus $\nabla_{h_{i}}|_{\mathcal{E}_{1}}=\nabla_{g_{1}}$ and in particular $h_{1}|_{\mathcal{E}_{1}}=g_{1}=h_{2}|_{\mathcal{E}_{1}}$ by Theorem 3.26 the uniqueness result for $\alpha^{n-1}$-stable sheaves (refer to Remark 3.1). Since $\mathcal{E}_{2}=E/\mathcal{E}_{1}$, it admits three $T$-adapted HYM metrics $g_{2}$ and $\widetilde{h_{i}}$ induced by $h_{i}$ and $h_{i}|_{\mathcal{E}_{1}}=g_{1}$ for $i=1,2$. Since $\mathcal{E}_{2}$ is $\alpha^{n-1}$-slope stable, we have that $\widetilde{h_{i}}=g_{2}$ by Theorem 3.26. Hence, inductively we obtain that
$h_{1}=h_{1}|_{\mathcal{E}_{1}}\oplus\cdots\oplus h_{1}|_{\mathcal{E}_{k}}=g_{1}\oplus\cdots\oplus g_{k}=h_{2}|_{\mathcal{E}_{1}}\oplus\cdots\oplus h_{2}|_{\mathcal{E}_{k}}=h_{2}.$
3.5 Asymptotic behavior of adapted HYM metrics toward $E_{nK}(\alpha)$
The proof of the estimates of $T$-adapted HYM metrics Theorem 3.13 and the uniqueness of $T$-adapted HYM metric Theorem 3.26, we obtain the following estimates.
Corollary 3.30.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$ such that $D=E_{nK}(\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. Let $T$ be an adapted closed positive $(1,1)$-current lying in $\alpha$. Assume that a holomorphic vector bundle $E$ admits a $T$-adapted HYM metric $h=h_{0}\Psi^{2}$ with $\det\Psi=1$. Then the following holds:
  1. $(1)$
    $\sup_{X\setminus D}|s_{D}\Psi|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty$.
  2. $(2)$
    We denote by $\Psi=e^{S}$ for some $h_{0}$-hermitian endomorphism $S$. Then $S$ is less singular that $\log|s_{D}|_{h_{D}}^{2}$ in the following sense: Let $\varepsilon_{k}>0$ be a sequence of constants so that $\varepsilon_{k}\searrow 0$. We denote by $D_{\varepsilon_{k}}$ the $\varepsilon_{k}$-neighborhood of $D$. Then
    $\inf_{k}\sup_{D_{\varepsilon_{k}}}\frac{|S|_{h_{0}}}{-\log|s_{D}|^{2}}=0.$
Proof.
Let $h=h_{0}\Psi^{2}$ be a $T$-adapted HYM metric on $E$. Then, by Theorem 3.28, the holomorphic vector bundle $E$ is $\alpha^{n-1}$-slope polystable. By Theorem 3.13 and Theorem 3.17, the endomorphism $\Psi_{\infty}$} in Theorem 3.13 gives a $T$-adapted HYM metric $h_{\infty}=h_{0}{\Psi_{\infty}}^{2}$ such that $\det{\Psi_{\infty}}=1$ and $\sup_{X\setminus D}|s_{D}{\Psi_{\infty}}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty$. By the uniqueness of $T$-adapted HYM metrics Theorem 3.26, we have $\Psi=a{\Psi_{\infty}}$. Since $\det\Psi=1=\det{\Psi_{\infty}}$, the constant $a$ equals to $1$: $\Psi={\Psi_{\infty}}$. Hence we obtain $(1)$.
Fix any integer $k\in\mathbb{Z}_{>0}$. Then $h^{\otimes k}=h_{0}^{\otimes k}\Psi^{\otimes k}$ gives a $T$-adapted HYM metric on $E^{\otimes k}$ with $\det\Psi^{\otimes k}=1$. Hence, there exists a constant $C_{k}>0$ such that
$\sup_{X\setminus D}\log(|s_{D}|_{h_{D}}|\Psi|^{k}_{h_{0}\otimes h_{0}^{*}})=\sup_{X\setminus D}\log|s_{D}\Psi^{\otimes k}|\leq C_{k}.$ (3.67)
We denote by $\Psi=e^{S}$ for some $h_{0}$-hermitian endomorphism $S$. Let $s_{1},\ldots,s_{r}$ be the eigenvalues of $S$ at $x\in X\setminus D$ and diagonalize $S$ at $x$ so that $S={\rm diag}(s_{1},\ldots,s_{r})$. Then, by (3.67) and by $|\Psi|^{2k}\geq e^{2ks_{1}}+\cdots+e^{2ks_{r}}$, we obtain
$C_{k}$ $\geq\log|s_{D}|^{2}+\log(e^{2ks_{1}}+\cdots+e^{2ks_{r}})$
$\geq\log|s_{D}|^{2}+\log(e^{2ks_{i}})$
$\geq\log|s_{D}|^{2}+2ks_{i}$ (3.68)
for any $i=1,\ldots,r$. By Theorem 3.14, the inverse morphism $\Phi=\Psi^{-1}$ also satisfies
$\sup_{X\setminus D}|s_{D}\Psi^{-1}|_{h_{D}\otimes h_{0}^{*}\otimes h_{0}}<\infty$. Then by applying the computation of (3.5) to $(\Psi^{\otimes k})^{-1}$, we obtain that for any $i=1,\ldots,r$,
$\log|s_{D}|^{2}-2ks_{i}\leq C_{k}.$ (3.69)
By (3.5) and (3.69), we obtain that $k|s_{i}|\leq C_{k}-\log|s_{D}|^{2}$. Hence we have
$k|S|_{h_{0}\otimes h_{0}^{*}}\leq r(C_{k}-\log|s_{D}|^{2}).$ (3.70)
Let us fix a sequence of constants $\varepsilon_{k}>0$ such that $\varepsilon_{k}\searrow 0$ and such that
$\frac{C_{k}}{-\log|s_{D}|^{2}}\leq 1$
on the $\varepsilon_{k}$-neighborhood $D_{\varepsilon_{k}}$ of $D$. Then by (3.70), we obtain
$\sup_{k}\sup_{D_{\varepsilon_{k}}}\frac{k|S|_{h_{0}\otimes h_{0}^{*}}}{-\log|s_{D}|^{2}}\leq 2r.$ (3.71)
Here we suppose that $\inf_{k}\sup_{D_{\varepsilon_{k}}}\frac{|S|_{h_{0}\otimes h_{0}^{*}}}{-\log|s_{D}|^{2}}\geq c>0$ for some constant $c>0$. Then
$\infty=\sup_{k}(kc)\leq\sup_{k}\sup_{D_{\varepsilon_{k}}}\frac{k|S|_{h_{0}\otimes h_{0}^{*}}}{-\log|s_{D}|^{2}},$
contradicting to (3.71). Hence we obtain
$\inf_{k}\sup_{D_{\varepsilon_{k}}}\frac{|S|_{h_{0}\otimes h_{0}^{*}}}{-\log|s_{D}|^{2}}=0.$
3.6 Semistable case
Definition 3.31.
[…]
$0=\mathcal{E}_{k+1}\subset\mathcal{E}_{k}\subset\mathcal{E}_{k-1}\subset\cdots\subset\mathcal{E}_{1}\subset\mathcal{E}_{0}=\mathcal{E}$
[…]
$Gr^{JH}(\mathcal{E}):=\bigoplus_{i=1}^{k+1}(\mathcal{E}_{i-1}/\mathcal{E}_{i}).$
[…]
Theorem 3.32.
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. Let $\mathcal{E}$ be a torsion free sheaf on $X$. Suppose that $\mathcal{E}$ is $\alpha^{n-1}$-slope semistable. Then the following holds.
  1. $(1)$
[…]
  2. $(2)$
[…]
Proof.
By Lemma 2.12, we can assume that $\mathcal{E}$ is locally free.
(1) Assume that $\mathcal{E}$ is $\alpha^{n-1}$-slope semistable, but not stable, Then, by [Jin25-2, Lemma 4.4], there exists a saturated subsheaf $\mathcal{E}_{1}\subset\mathcal{E}$ such that
  • (a)
    $\mu_{\alpha}(\mathcal{E}_{1})=\mu_{\alpha}(\mathcal{E})$ and $0<{\rm{rk}}\mathcal{E}_{1}<{\rm{rk}}\mathcal{E}$.
  • (b)
    For any saturated subsheaf $\mathcal{F}\subset\mathcal{E}$ satisfying the two above conditions, then $\mathcal{E}_{1}\not\subset\mathcal{F}$ or $\mathcal{E}_{1}=\mathcal{F}$.
Then the rest of the proof of (1) is standard as follows. By (a) above, we have $\mu_{\alpha}(\mathcal{E}/\mathcal{E}_{1})=\mu_{\alpha}(\mathcal{E})$. We prove that $\mathcal{E}/\mathcal{E}_{1}$ is $\alpha^{n-1}$-slope stable. Let $\mathcal{F}\subset\mathcal{E}/\mathcal{E}_{1}$ be a saturated subsheaf with $0<{\rm{rk}}\mathcal{F}<{\rm{rk}}(\mathcal{E}/\mathcal{E}_{1})$. We consider the following short exact sequences
[…]
Since $\mathcal{E}$ is $\alpha^{n-1}$-slope semistable, we have $\mu_{\alpha}(\mathcal{E}_{1}+\mathcal{F})\leq\mu_{\alpha}(\mathcal{E})$. Then by the exact sequence in the second line, we have
$0$ $={\rm{rk}}\mathcal{F}(\mu_{\alpha}(\mathcal{E}_{1}+\mathcal{F})-\mu_{\alpha}(\mathcal{F}))+{\rm{rk}}\mathcal{E}_{1}(\mu_{\alpha}(\mathcal{E}_{1}+\mathcal{F})-\mu_{\alpha}(\mathcal{E}_{1}))$
$\leq{\rm{rk}}\mathcal{F}(\mu_{\alpha}(\mathcal{E})-\mu_{\alpha}(\mathcal{F}))+{\rm{rk}}\mathcal{E}_{1}(\mu_{\alpha}(\mathcal{E})-\mu_{\alpha}(\mathcal{E}_{1}))$
$={\rm{rk}}\mathcal{F}(\mu_{\alpha}(\mathcal{E}/\mathcal{E}_{1})-\mu_{\alpha}(\mathcal{F})).$
Hence we have $\mu_{\alpha}(\mathcal{F})\leq\mu_{\alpha}(\mathcal{E}/\mathcal{E}_{1})$. If $\mu_{\alpha}(\mathcal{F})=\mu_{\alpha}(\mathcal{E}/\mathcal{E}_{1})$, then the first line above shows
${\rm{rk}}\mathcal{F}(\mu_{\alpha}(\mathcal{E}_{1}+\mathcal{F})-\mu_{\alpha}(\mathcal{E}_{1}))=0.$
[…]
Finally we formulate conjectures related to the above Theorem 3.32. It is a generalization of [Sib15]. Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class on $X$. We assume that $D=E_{nK}(\alpha)$ is a snc divisor. We denote by $s_{D}$ a defining section of $D$ and $h_{D}$ a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$. Let $T$ be a adapted closed positive $(1,1)$-current in $\alpha$. Let $E$ be a holomorphic vector bundle with a smooth hermitian metric $h_{0}$. Then we consider the space of $T$-adapted hermitian metrics on $E$ denoted by $\mathcal{T}_{h_{0}}$. That is, $\mathcal{T}_{h_{0}}$ consists of a smooth hermitian metric $h=h_{0}\Psi^{2}$ on $E|_{X\setminus D}$ such that $\|\Lambda_{T}F_{h}\|_{L^{\infty}(X\setminus D,h_{0})}<\infty$, $\int_{X\setminus D}|F_{h}|_{h,T}T^{n}<\infty$, $\sup_{X\setminus D}|s_{D}\Psi|_{h_{D}\otimes h_{0}}<\infty$ and $\int_{X\setminus D}|s_{D}^{m+1}\nabla_{h_{0}}\Psi|^{2}_{h_{D}^{m+1}\otimes h_{0}\otimes h_{0}^{*},T}T^{n}<\infty$. We give a norm of $h=h_{0}\Psi^{2}\in\mathcal{T}_{h_{0}}$ by
$\|h\|:=\sup_{X\setminus D}|s_{D}\Psi|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}$
$\hskip 11.38109pt+\left(\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi|_{h_{D}\otimes h_{0}\otimes h_{0}^{*},T}T^{n}\right)^{1/2}+\int_{X\setminus D}|s_{D}^{m+1}\Delta_{T}\Psi|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}T^{n}.$
Then we can formulate the following conjecture.
Conjecture 3.33.
  1. $(1)$
    For any holomorphic vector bundle $E$, there exists $\{h_{t}\}_{t\in[0,\infty)}$$\subset\mathcal{T}_{h_{0}}$ such that
    $h_{t}^{-1}\frac{d}{dt}h_{t}=-(\sqrt{-1}\Lambda_{T}F_{h}-\lambda{\rm{id}})$
    for $t\in[0,\infty)$.
  2. $(2)$
[…]